### v1

# Range Membership Inference Attacks

Abstract
Machine learning models can leak private information about their training data, but the standard methods to measure this risk, based on membership inference attacks (MIAs), have a major limitation. They only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping data revealing the same private information. To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy). We formulate the RaMIAs game and design a principled statistical test for its complex hypotheses. We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language. RaMIA paves the way for a more comprehensive and meaningful privacy auditing of machine learning algorithms.
1 Introduction
Machine learning models are prone to training data memorization [14, 15, 27, 41, 24]. It is also a known fact that the outstanding predictive performance of machine learning models on long-tailed data distributions often comes at the expense of blatant memorization of certain data points [15, 3, 30, 16]. In simple words, memorization is the phenomenon that models behave differently on training points, compared to other points. The memorization can lead to significant privacy risks as adversaries can infer private information about the training data from only having black box access to models.
To quantify the privacy risk of machine learning models, a privacy notion needs to be fixed first. The reigning privacy notion is defined by membership information. Membership information of a data point is binary, but this single bit of information carries huge privacy implications. Being able to infer membership information opens up the possibility of conducting data reconstruction attack [36, 17, 4, 29], where the reconstruction attack inspects the membership of plausible data points to recover the training set. The de-facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [39], where an adversary aims to predict whether a given query data is part of the training set of the target model. The more powerful the membership inference attack is, the higher the privacy risk the target model bears.
Membership inference attacks provide a lower bound of the true privacy risk of a model, so improving the attack performance also means tightening the bound of privacy risk estimation. So far, the community has put much effort into improving the power of membership inference attacks by crafting better membership signals and constructing better statistical tests [38, 39, 35, 44, 5, 47]. While these have been useful for the betterment of privacy auditing, they have ignored the fundamental drawback of membership inference attacks as a practical privacy auditing tool, i.e., MIAs assume it is a privacy concern only if the adversary can identify the exact, full version of the training data. However, if the adversary can identify data points that are similar enough to the training points, it should also be treated as a significant privacy risk, because those points can contain similar levels of private information. For example, two Alice’s photos taken from slightly different angles, or with a different background would contain similar private information about Alice’s face or location. Similar to images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13]. This oversight means private information leakage beyond exact matches of training data is ignored and current privacy auditing tools might produce overly optimistic results.
Besides, focusing on exact membership inference attacks renders them incapable of handling queries with missing values. This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features would carry a similar level of private information as the full data records. Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in the hospital training set, there is no need to identify the rest of the features because the adversary is already able to pinpoint who has HIV from the given subset of key features. Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features that often contain quasi-identifiers, which has been studied extensively in the failure of k-anonymity [37], where the attacker is able to reconstruct and identify training data with certain features removed. However, if we make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member," since the chance of us coming up with the right values is so slimSupposeIf we use property inference [40, 1, 8, 9, 45, 20] to help fill in the missing features, the imputed data may be have inflated membership score even if they are non-members, leading to a higher false positive rate. Therefore, existing frameworks struggle with quantifying privacy leakage with missing features.
We argue that privacy quantification should not be point-based, because a small neighborhood around training points also contains similar private information. Hence, in this paper, we are proposing a new attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy. Instead of using point queries and testing for exact matches, range membership inference attacks use range queries that cover a set of points. The goal of range membership inference attacks is to infer if the given range query contains any training point.
Range membership inference attacks extend the formulation of membership inference attacks. We adapt the original inference game formulation to reflect the change to range queries in RaMIAs. This extended formulation produces composite hypotheses in the likelihood ratio tests, which are the standard and best attack techniques in MIAs [38, 44, 5, 47]. Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests and Bayes factors. We show that RaMIAs can provide a more comprehensive notion of privacy by detecting private information leakage from the vicinity of training data when MIAs underestimate such privacy risk. Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 0.20 if we test image classifiers with horizontally flipped images (Figure 2(c)). RaMIA, implemented with our simple attack strategy (Sec 4), supersedes MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
In this paper, we emphasize the motivation and formulation of our newly proposed attack framework, RaMIA. As a proof-of-concept, we experiment RaMIA with a simple attack strategy on tabular, image and text datasets, where RaMIA unanimously outperforms MIA. Additionally, our attack can also be potentially used in the pioneering membership inference and data extraction attacks on generative models [4, 43, 7], where the current evaluation requires finding the closest training image for all candidate data and computing their distances. By setting the distance function as the range function and conducting RaMIAs, we can evaluate the attacks more systematically.
(a) MIAScore=2.99
(b) MIAScore=2e-15
(c) MIAScore=1.8 and 1.0
Figure 1: Examples of similar data with drastically different MIA scores. MIA would classify one of them as member, but the other as non-member.
2 Preliminaries
2.1 Membership inference attacks
The membership inference attack (MIA) [39] is a type of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set. Mathematically, given a model $f$ and a query point $\mathbf{x}$, the MIA aims to output 1 if $\mathbf{x}$ is a training point, and 0 otherwise. There are various methods to construct and conduct the attack, and it is still an active research direction that sees more powerful attacks being developed. Shokri et al. [39] use a shadow model based approach where shadow models are trained on known training sets in similar ways to the target model. Confidence values of the training and test data on the shadow models are computed, which are then used as benchmarks in testing. However, the high cost and strong assumption of knowing the target model’s training details make the attack often infeasible. Yeom et al. [45] use model loss as a signal and threshold it, scraping the need for shadow models. Then MIA is formulated as an inference game (See Sec 3.1.1). Researchers turn to the principled approach to solve the game via likelihood ratio tests [38, 5, 44, 47]. Carlini et al. [5] and Ye et al. [44] propose reference-model based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio. To further boost the attack power, Zarifzadeh et al. [47] assume the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
Recent attacks [4, 47] further boost the attack performance on image data by augmenting the test queries with train-time augmentations. This assumes that the attacker knows the exact train-test augmentations in advance, and is able to sample from them. Augmenting training images with non train-time augmentations is not considered for valid reasons: those augmented images would be non-members in the current privacy notion.
2.2 Range queries
If we make a connection to the field of databases, a membership inference attack operates on point queries or exact match queries. That is, each query to the membership inference attack only contains one data point and the attack only concerns if this very point is in the training set. On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the "range". The most fundamental difference to point query is that the retrieved result often contains multiple data points instead of a single one. Our proposed attack, the range membership inference attack, operates with range queries.
3 From MIA to RaMIA
Membership inference attacks are often formulated as an inference game [45, 21, 44, 5, 47] between a challenger and an adversary. In this section, we will walk through how we come up with RaMIA from MIA.
3.1 Membership inference attacks
In membership inference attacks, the goal is to identify if a given point is part of the training set.
3.1.1 Membership inference game
Definition 1
(Membership Inference Game [44, 45]) Let $\pi$ be the data distribution, and let $\mathcal{T}$ be the training algorithm.
  1. 1.
    The challenger samples a training dataset $D\overset{s_{D}}{\longleftarrow}\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
  2. 2.
    The challenger samples a data record $z_{0}\overset{s_{z_{0}}}{\longleftarrow}\pi$ from the data distribution, and a training data record $z_{1}\overset{s_{z_{1}}}{\longleftarrow}D$.
  3. 3.
    The challenger flips a fair coin to get the bit $b\in\{0,1\}$, and sends the target model $\theta$ and data record $z_{b}$ to the adversary.
  4. 4.
    The adversary gets access to the data distribution $\pi$ and access to the target model, and outputs a bit $\hat{b}\longleftarrow\mathcal{A}(\theta,z_{b})$.
  5. 5.
    If $\hat{b}=b$, output 1 (success). Otherwise, output 0.
3.1.2 Evaluation of MIA
Evaluation is done with a set of training and test points. True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all possible threshold values. By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC). A clueless adversary who can only randomly guess the membership labels will get an AUC of 0.5. For stronger adversaries, they predict membership more correctly at each error level. Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
3.1.3 Intrinsic limitation of MIA as a Privacy Auditing Framework
MIAs are intrinsically incapable of identifying points close to training points, regardless of how similar they are, because these points are, by definition, non-members in the scope of MIAs. Hence, there is a huge space of points that contain private information but are deemed non-members in the current privacy auditing framework. In this way, MIAs as privacy auditing tools become bad when the queries move away from the original data. Figure 2 shows the MIAs under-perform on non-original data. This inspires our formulation of RaMIA, where these points will be classified as "members" for better and more comprehensive privacy auditing.
(a) Purchase-100
(b) CelebA
(c) CIFAR-10
(d) AG News
Figure 2: MIA performance gets worse when the query points become further away from the training points. We define points different from training points but carry similar information as members. In 2(a), $m$ is the number of missing values. The query is constructed by filling in them with the most likely values. In 2(b), the point query changes to photos of the same identity who has at least one photo in the training set. In 2(c), the queries are vertically flipped images. In 2(d), $d$ is the Hamming distance to original sentences.
3.2 Range membership inference attack
In range membership inference attacks, the goal is to identify if a given range contains any training point.
3.2.1 Range membership inference game
Here we define our range membership inference game, modified from the above formulation.
Definition 2
(Range Membership Inference Game) Let $\pi$ be the data distribution, and let $\mathcal{T}$ be the training algorithm.
  1. 1.
    The challenger samples a training dataset $D\overset{s_{D}}{\longleftarrow}\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
  2. 2.
    The challenger samples a data record $z_{0}\overset{s_{z_{0}}}{\longleftarrow}\pi$ from the data distribution, and a training data record $z_{1}\overset{s_{z_{1}}}{\longleftarrow}D$.
  3. 3.
    The challenger flips a fair coin to get the bit $b\in\{0,1\}$. If $b=1$, the challenger samples a range $\mathcal{R}_{1}$ containing at least one training point. Otherwise, challenger samples a range $\mathcal{R}_{0}$ containing no training points.
  4. 4.
    The challenger sends the target model $\theta$ and the range $\mathcal{R}_{b}$ to the adversary.
  5. 5.
    The adversary gets access to the data distribution $\pi$ and access to the target model, and outputs a bit $\hat{b}\longleftarrow\mathcal{A}(\theta,\mathcal{R}_{b})$.
  6. 6.
    If $\hat{b}=b$, output 1 (success). Otherwise, output 0.
The main difference between the two games is that the adversary now receives a range query (Step 4 in Def 2) instead of a point query (Step 3 in Def 1). We assume that the adversary is able to sample within the range. This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [39, 44, 47]. Given a range query and a sampler of $\pi$, it is not difficult to sample within the range.
What is a range
A range can be defined by a center, which is a point, a radius representing the size of the range, and a distance function which the radius is defined with. We refer to the center as the query center, the radius as the range size, and the distance function as the range function in this paper. One way to visualize a range is to imagine a unit $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions. Our framework can cater to any arbitrary range function. It can be spatial based (e.g. $l_{p}$ distances), transformation based (e.g. geometric transformations) and semantic based (e.g. owner/main features of the data). In the experiment section, we will present results with all of these types of range functions. Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based.
How to construct a range
In Step 3 of the range membership inference game, the details of how the challenger samples the ranges are intentionally omitted. This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both in- and out-ranges. The details of how we construct the in- and out-ranges for our experiments are elaborated in Appendix B.
3.3 Evaluation of RaMIA
Similarly, we evaluate RaMIA with AUCs. However, the notion of true positives and false positives are different from those defined in MIA, as both are defined on the range level. To avoid confusion, we call them Range TPR and Range FPR, which means a range is correctly/wrongly predicted to contain at least one training point.
4 Range membership inference attacks
4.1 Composite hypothesis testing
Similar to likelihood ratio tests for membership inference games A.1, we can also construct two hypotheses for range membership inference game (Def 2):
[…]
$\forall z\in\mathcal{R}_{b}:z\not\in D.$
[…]
$\exists z\in\mathcal{R}_{b}\text{ s.t. }z\in D.$
[…]
To make full use of the Bayes Factor, we need to know the prior distribution, which is unrealistic. On the other hand, taking the max seems to be a more intuitive approach for range membership inference attacks, because it provides a two-step solution: search and test. Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points. However, this assumes that we can reliably find the max values in a given range. Since most ranges are large data subspaces, it is very challenging to find the optimal points within the large space. Even if the search space can be navigated, any search algorithm is likely to return local maxima. Hence, a robust way is to aggregate the top samples. However, membership inference attacks are known to be unreliable on out-of-distribution (OOD) data [47] and can assign them high scores. When the sampling space contains only these data as opposed to real and in-distribution (ID) data, the maximum might not be anything close to true training data, increasing the FPR and lowering the AUC as a result.
In this paper, we adopt a simple attack strategy. Our solution to this is based on the type of data in the sampling space. If the data are all naturally ID, we can combine Bayes Factor and GLRT by taking the average likelihood of the top samples to reduce the influence of the randomness in the sampling process. On the other hand, if the adversary can only synthesize data within the range, the top samples are highly likely to be OOD data with high membership scores. Hence, in this case, we want to remove those points from the equation. Since the presence of training points intuitively raises the average membership score of ID points nearby, compared to having no training points at all, we average the membership scores of the remaining samples. Unifying these two strategies gives us the following:
$\mathbb{P}\left(\theta|H_{1}\right)=\text{TrimmedAvg}(S,q_{s},q_{e};\mathbb{P})=\text{Avg}_{x\not\in[q_{s},q_{e}]\text{-th quantiles}}\mathbb{P}(\theta|x\in D),$ (1)
where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles where we want to remove to compute our robust statistics that are one-sided trimmed means. If the sampling space is filled with synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points. $q_{s}$ is a hyperparameter that decreases (trim more) as the quality of sampled points gets worse. On the other hand, if the sampling space consists of real points, we set $q_{s}=0$ to remove the smallest points in our aggregation. $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available. Note that the optimal hyperparameters may differ across different membership signals (e.g. loss values, LiRA scores), as they exploit different vulnerabilities and expose different training points. However, for fixed model architectures, range functions, data distributions and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [47]. Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters.
4.2 Range membership inference attack as a framework
The range membership inference attack is a new inference attack framework, not a particular attack algorithm. There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 1. The sampler $\text{Sampler}(\mathcal{R}):S\rightarrow X$ returns samples within the given range. The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\mathbb{P}(\theta|x)$. A number of existing attack algorithms can be plugged in. Similar to MIAs, the key to using RaMIA as a privacy auditing tool is to compute the range membership score. Our framework can adopt any existing membership scoring function $\text{MIA}(x)$ to compute $\text{RaMIA}(\mathcal{R})$. Below, we outline the attack with our attack strategy described above:
Algorithm 1 Computing range membership scores
1:Input range $\mathcal{R}$, sampler $\text{Sample}(\cdot)$, target model $\theta$, membership scoring function $\text{MIA}(\cdot)$.
[…]
3:if samples are real and ID then
4:     Set $q_{s}=0$, and set $q_{e}$ by sweeping on reference models;
5:else
6:     Set $q_{e}=100$, and set $q_{s}$ by sweeping on reference models.
7:end if
8:$\text{RaMIA}(\mathcal{R};\theta)=\text{TrimmedAvg}(S,q_{s},q_{e};\text{MIA})$
5 Experiments
Since the purpose of this paper is to introduce a new concept and framework, the goal of the experiments section is to provide a proof-of-concept. We experiment on the commonly used Purchase-100 [39], CelebA [28], CIFAR-10 [25] and AG News [48] datasets. Details of how we split the dataset, train models, construct ranges, and obtain samples are explained in Appendix B. We compare RaMIA with MIA (Sec 5.1) in scenarios MIA under-performs (Depicted in Figure 2). Both attacks are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [47], with three reference models trained in the same way as Carlini et al. [5] and Zarifzadeh et al. [47]. The respective queries are outlined in Table 1, and the definitions of members under each attack framework are explained in Table 2. In both tables, $x$ represents original data in datasets, while $x^{\prime}$s are either data with missing values or modified data from $x$. The reason that we do not test our attacks by taking ranges centered at original data $x$ is that the chance of the attack data being exactly the same as the training data is extremely low without sufficient prior knowledge. It is more realistic that similar data are being queried.
Table 1: Range queries and point queries used in our experiments for RaMIA and MIA respectively.
Dataset Range query Point query
Purchase-100 possible data records given the incomplete data $x^{\prime}$ mode imputed $x^{\prime}$
CelebA photos featuring the same person as photo $x^{\prime}$ photo $x^{\prime}$
CIFAR-10 transformed versions of image $x^{\prime}$ image $x^{\prime}$
AG News sentences that are of Hamming distance 8 to sentence $x^{\prime}$ sentence $x^{\prime}$
Table 2: Definitions of range and point members corresponding to the attack queries in Table 1
Dataset Range member if there is at least (Point) member if
Purchase-100 one training point matches with $x^{\prime}$ on all unmasked columns $x^{\prime}_{\text{impu}}$ is member
CelebA one training image featuring the same person as $x^{\prime}$ $x^{\prime}$ is member
CIFAR-10 one version of image $x^{\prime}$ in the training set $x^{\prime}$ is member
AG News one training sentence within Hamming distance 8 to $x^{\prime}$ $x^{\prime}$ is member
On Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$. On CIFAR-10, we apply up to 15 distinct transforms, and set $q_{e}=100,q_{s}=40$. On AG News, we construct 50 sentences within each range, and set $q_{e}=100,q_{s}=20$. On CelebA, each celebrity has a different number of images in the sampling space, ranging from 1 to 18. Since it is hard to standardize the sample size for all ranges, we take all of them. We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
5.1 RaMIAs quantify privacy risks more comprehensively than MIAs
As we have explained before, data points that are close enough to the training data are out of the scope of membership inference attacks. We observe from Figure 3 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested. We want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes. On Purchase-100, there are a total of 1024 candidates, and we take less than 20% of them. On AG News, there are millions of sentences within a distance of 8. 50 sentences are too little to meaningfully cover anything in the space. But yet limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is flawed and needs a better framework. Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 4.TPRs at small FPRs are in Table 3
(a) Purchase-100 (m=10)
(b) CelebA
(c) CIFAR-10
(d) AG News (d=5)
Figure 3: RaMIAs perform better than MIAs on points that are close to original points but not exactly the same.
5.2 Factors affecting RaMIA performance
Training data density in the range
Due to the nature of the sampling-based approach, the chance of our attack set containing a true training point scales linearly with the density of training points in the range. If we keep the sample size constant, increasing the range without including more training points in the range hurts the attack performance because the chance of the attack set including any training point gets diluted. On the other hand, if we increase the training point density, which is equivalent to increasing the probability the attacker samples a true training point, the attack performance gets boosted. Figure 4(b) shows that the performance of RaMIA increases when the range becomes larger in the CIFAR-10 experiment. Recall that the range function in CIFAR-10 is based on image augmentation methods. Increasing the range means the attacker applies more distinct augmentation methods to obtain transformed images. This increases the chance of the attacker obtaining one of the transformed versions of training images seen by the model during training, thus leading to better attack performance. In Figure 3(b), we conducted the attack assuming the attacker cannot sample any true training images. As a sanity check, we relax this assumption, and Figure 4(a) shows that RaMIA performs monotonically better when the density of training images increases from 0% to 50%, when the number of samples is constant.
(a) RaMIA on CelebA gets better when the training points available for sampling increases. The percentages are the density of training points in the distribution that the attack samples from. (Sanity Check)
(b) RaMIA performs better on CIFAR-10 when the range size increases. The size is equal to the number of distinct transformation applied to images.
Figure 4: Attack performance increases in tandem with training point density.
Susceptibility to MIAs and RaMIAs is correlated
Ranges containing training points that are susceptible to MIAs are also more susceptible to RaMIAs. Researchers have previously discovered that machine learning models memorize duplicate data more [26, 6]. In our CelebA dataset, each celebrity has different numbers of photos in the training set, which can be thought as that each identity has different levels of duplication in the training set. Similar to the insights from MIAs, we also observe that identities that have more training images, i.e. higher duplication rate, are more susceptible to RaMIA. Figure 8 shows the relationship between the percentile each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate. Generally speaking, identities that have more training photos are more prone to RaMIAs. Similarly correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 7.
5.3 Mismatched training and attack data hurts attack performance
Figure 2(c) shows that MIA underestimates the privacy risk when the augmentation used in training and attacking differs. This rings a bell as many people audit the privacy risk of image classifiers with original images, when the classifiers are often trained with a composition of augmentations. Many transformations, such as color jittering and affine transformations, always produce different final images. Other commonly used augmentation methods, such as random cropping, introduce more randomness to the pipeline. Hence, it is almost certain that the original images are never seen by the model. Therefore, we should use RaMIA for a better auditing result (Figure 3(c)).
Difference to existing augmentation-based MIAs
Existing attacks [4, 47] also use augmented queries in the attack, but with a different rationale and assumption of the attacker’s knowledge. Since they adopt the existing privacy notion based on point queries, only (augmented) images seen by the model in the training stage are considered as members. Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy of notion. In RaMIA, the set of augmentations is given by the challenger (Def 2), which can contain augmentations not used in training, but considered as privacy leaking. Using the aggregation method in [4] will hurt the attack performance if non-training augmentations are used. However, RaMIA is designed to be robust in this scenario (Fig 3(c)).
6 Conclusion
In this paper, we argue that membership inference attacks are only useful as a privacy audit tool when querying exact copies of training and test data. Moving the query to similar points causes a drastic decrease in performance, rendering MIAs less useful. We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs. and cover the failure cases of MIAs by checking if a given range contains a training point. We introduce RaMIA as an attack framework that can be implemented with any existing MIA algorithm. We show that it can provide better privacy auditing with very few samples taken randomly. We hope our work can make more privacy researchers and practitioners aware of the shortcomings of MIAs, and shift their attention to RaMIAs. As it is the first paper that brings up this new framework, there is room for improvement in specific attack algorithms. For example, a better sampling process will surely increase the gap between RaMIA and MIA. Nevertheless, we have shown our framework is sensible and useful. In future work, we hope to design more powerful RaMIA strategies that are robust to the change of membership signals and datasets, especially on LLMs where we believe our privacy notion is extremely relevant.
Appendix A Attack Details
A.1 (Simple) Hypothesis testing
The standard way to tackle the inference game (Def 1) is to apply statistical hypothesis tests [44, 5]:
$H_{0}$ $:\text{The given }z\text{ is not a training point }(b=0).$
$H_{1}$ $:\text{The given }z\text{ is a training point }(b=1).$
The likelihood ratio test (LRT) is then conducted
$\frac{\mathbb{P}\left(\theta|H_{1}\right)}{\mathbb{P}\left(\theta|H_{0}\right)}$ (2)
This is usually called "simple" hypothesis testing because each $H$ contains a single hypothesis.
A.2 Attack algorithms
In this section, we explain the details of the membership inference attack algorithms used in our experiments.
LOSS
LOSS [45] computes loss values as a proxy of membership score on given points: $\text{MIA}(x;\theta)=\mathit{l}(x;\theta)$. To compute the likelihood, an easy way is to take the exponential of the negative of the loss $P=\exp^{-\mathit{l}}$.
RMIA
RMIA [47] computes membership score by applying chain rule: $\mathbb{P}(\theta|x)=\frac{\mathbb{P}(x|\theta)\mathbb{P}(\theta)}{\mathbb{P}(x)}$. The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\mathbb{P}_{z\in Z}[\frac{\mathbb{P}(\theta|x)}{\mathbb{P}(\theta|z)}\leq\gamma]$, where the term $\mathbb{P}(\theta)$ will cancel out with each other. The normalizing constant $\mathbb{P}(x)$ is computed with reference models: $\mathbb{P}(x)=0.5\mathbb{E}_{\theta_{\text{in}}}\mathbb{P}(x|\theta_{\text{in}})+0.5\mathbb{E}_{\theta_{\text{out}}}\mathbb{P}(x|\theta_{\text{out}})$. In its offline version, the in models are unavailable. In this case, the former probabilities is approximated by the latter term $\mathbb{P}_{\text{in}}=a\mathbb{P}_{\text{out}}+(1-a)$. The hyperparameter $\alpha$ is chosen based on the reference models. Specifically, one reference model is chosen as the temporary target model, and the rest are used to attack it. The value of $\alpha$ is chosen to be the best performing value under this setting, obtained via a simple sweeping. In our experiment, we use the offline attack only. The $\alpha$ values for Purchase-100 and CIFAR-10 are taken from [47]. For CelebA, we set it to be 0.33. For AG News, we set it to be 1.0.
Appendix B Setup Details
On each of the dataset, we train four models on half of the dataset in the same way described by Carlini et al. [5], Zarifzadeh et al. [47]. We will describe the details of the datasets below. We have checked with their licenses with our best effort, and confirm their terms of use are respected.
B.1 Tabular data: Purchase-100
Dataset
Purchase-100 [39] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data. This dataset was first curated by Shokri et al. [39] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product. The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
Models
We train five-layer multi-layer perceptron (MLP) models in PyTorch [33] on half of the entire dataset. The hidden layers are of sizes $[512,256,128,64]$. All models achieve a test accuracy of $83\%$.
Construction of ranges
We simulate the scenario where the attacker has incomplete data (data with missing values). For all training and test data records, we randomly mask k columns. Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary. We then check if each range constructed from test data points include any training point. If so, they are re-labelled as "in-ranges".
Sampling within ranges
Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns. The parameter of the sampler is computed by taking the average value of each column. Because of the nature of this dataset, our sampled data can be regarded as in-distribution. We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
B.2 Image data I: CelebA
Dataset
CelebA [28], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features. We construct the members set by only including photos of celebrities with identity number smaller than 5090. The rest are used to construct the non-members set. For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
Model
We train four-layer convolutional neural networks (CNNs) in PyTorch [33] on the training set to predict the facial attributes of any given photo. Our target model has a test accuracy of $87\%$.
Construction of ranges
The range function here is a semantic one that is based on the identity of the face image. For example, a range query can be "all Alice’s photos". Since the identities in the training and non-members set are disjoint, it is easy to construct in- and out-ranges based on the distribution of identities in the two sets.
Sampling within ranges
For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
B.3 Image data II: CIFAR-10
Dataset
CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
Models
We train WideResNets-28-2 [46] with JAX [2] on half of the training set of CIFAR-10 using the code from [5], with and without image augmentations. Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$. The train time augmentation is the composition of random flipping, cropping and random hue.
Construction of ranges
The range function here is different types of image augmentations, which are geometric transformations. An example of a range query is "all transformed version of image $X$". For each training and test image, a range is constructed by applying different transformations.
Sampling within ranges
For each range query, we independently apply 15 image augmentations on the query center. The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
B.4 Textual data: AG News
Dataset
We use AG News dataset [48], which is a news collection with four categories of news. Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset. There are 120,000 sentences in its training set.
Model
We took pretrained GPT-2 [34] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library. The finetuning is done for 4 epochs. Our target model achieves a perplexity of 1.39 on the test set of AG News.
Construction of ranges
The range function here is word-level Hamming distance, which can be thought as the edit distance measured on word level that only allows word substitution. An example of a range query is "all sentences within Hamming distance $d$ to sentence $x$". To construct in- and out-ranges, we just need to specify the max Hamming distance and the starting sentence. We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pre-trained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences. A Hamming distance is then specified with each starting sentence to form a range.
Sampling within ranges
We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
Appendix C Implementation details
For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001. For WideResnets, we use the training code from [4]. On AG News, the models are trained for 4 epochs. On other datasets, they are trained for 100 epochs.
All training are done on two Nvidia RTX 3090 GPUs. Training on AG News takes about 1 hour per epoch. Training other models takes less than one hour each.
Appendix D RaMIA on redacted data
Many large language models (LLMs) are trained with sensitive textual data. Some of the data with sensitive information redacted might be public available. Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM. Accurately identify the redacted sentences paves the way for reconstructing them as a follow-up attack. Figure 5 shows the results. In this experiment, we use spaCy [18] to mask peoples’ names to simulate masking of personally identifiable information (PII). We then generate 10 possible sentences for each masked sentence using BERT, and conduct RaMIA. The MIA performance is the average attack performance over all 10 possible sentences.
Figure 5: RaMIA on subset of AG News where names are redacted.
Appendix E Extra results
In this section, we put extra experiment results.
Table 3: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$. MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them. Standard deviation over random sampling iterations are reported, except for CelebA, where we use all available candidates.
Purchase-100 CIFAR-10 CelebA AG News
TPR@FPR(%) 1% 0.1% 1% 0.1% 1% 0.1% 1% 0.1%
MIA
LOSS 0 0 0.15 0 1.86 0.31 $80.0$
RMIA 2.18 0.37 2.40 0.21 1.69 0.19 $0.30$ 00
RaMIA
LOSS $0\pm 0$ $0\pm 0$ $1.17\pm 0.06$ $0.05\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $3.59\pm 0.11$ $0.80\pm 0.02$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
Table 4: Improvement in AUCs after switching from MIA to RaMIA across multiplue iterations of random sampling. We do not random sample but use all available attack images in CelebA.
Purchase-100 CIFAR-10 CelebA AG News
$\Delta\text{AUC}$ $2.62\pm 0.04$ $4.12\pm 0.06$ $5.4$ $1.20\pm 0.2$
Figure 6: RaMIA gets better when the range size increases on CelebA, even if we assume the attacker cannot sample any training image.
(a) Purchase-100
(b) CIFAR-10
(c) AG News
Figure 7: Correlation between the percentile of RaMIA and MIA scores of members among non-members. The larger the percentile is, the more non-members the member dominates, and the more likely for it to be classified as IN. The Pearson correlation coefficients are provided for each plot. It shows that the vulnerability to MIA is correlated with that to RaMIA.
Figure 8: Number of photos from the same celebrity in the training set affects the identifiability of the range.

### v2

@@
 To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy).
-We formulate the RaMIAs game and design a principled statistical test for its complex hypotheses.
+We formulate the RaMIAs game and design a principled statistical test for its composite hypotheses.
 We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language.
@@
 The memorization can lead to significant privacy risks as adversaries can infer private information about the training data from only having black box access to models.
-To quantify the privacy risk of machine learning models, a privacy notion needs to be fixed first.
+To quantify the privacy risk of machine learning models, a privacy notion needs to be agreed upon first.
 The reigning privacy notion is defined by membership information.
@@
 For example, two Alice’s photos taken from slightly different angles, or with a different background would contain similar private information about Alice’s face or location.
-Similar to images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13].
-This oversight means private information leakage beyond exact matches of training data is ignored and current privacy auditing tools might produce overly optimistic results.
+Similar to geometric transformations in images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13].
+This oversight means private information leakage beyond exact matches of training data is currently ignored and the contemporary privacy auditing tools might produce overly optimistic results.
 Besides, focusing on exact membership inference attacks renders them incapable of handling queries with missing values.
-This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features would carry a similar level of private information as the full data records.
-Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in the hospital training set, there is no need to identify the rest of the features because the adversary is already able to pinpoint who has HIV from the given subset of key features.
-Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features that often contain quasi-identifiers, which has been studied extensively in the failure of k-anonymity [37], where the attacker is able to reconstruct and identify training data with certain features removed.
-However, if we make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member," since the chance of us coming up with the right values is so slimSupposeIf we use property inference [40, 1, 8, 9, 45, 20] to help fill in the missing features, the imputed data may be have inflated membership score even if they are non-members, leading to a higher false positive rate.
-Therefore, existing frameworks struggle with quantifying privacy leakage with missing features.
-We argue that privacy quantification should not be point-based, because a small neighborhood around training points also contains similar private information.
-Hence, in this paper, we are proposing a new attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy.
+This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features could carry a similar level of private information as the full data records.
+Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in an HIV patients dataset, there is no need to identify or collect the exact values of the rest of the features, because the adversary is already able to pinpoint who has HIV from the given subset of key features.
+Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features from the subset of known features that often contain quasi-identifiers, which has been extensively studied in the failure of k-anonymity as a privacy-preserving measure [38], where the attacker can reconstruct and identify training data with certain features masked.
+With existing inference attacks, there are two types of approaches to work with incomplete data.
+The first approach is to impute the missing features before performing membership inference attacks.
+However, if we simply make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member,” since the chance of us coming up with the right values is extremely slim.
+If we use attribute inference [20] to help fill in the missing features, the imputed data may have inflated membership scores even if they are non-members, leading to a higher false positive rate.
+Therefore, existing MIA-based privacy auditing frameworks does not provide a natural solution to quantifying privacy leakage with missing features.
+The second approach that works with incomplete data is property inference [41, 1, 8, 9, 46], but it aims to extract statistical information about the features instead of testing membership information of data records.
+Hence, none of the existing inference attack frameworks can reliably quantify privacy leakage with incomplete data entries.
+We argue that privacy quantification should not be point-based, because a small neighborhood around training points, defined based on a distance function out of the semantics of privacy, also contains similar private information.
+Hence, in this paper, we are proposing a new inference attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy.
 Instead of using point queries and testing for exact matches, range membership inference attacks use range queries that cover a set of points.
 The goal of range membership inference attacks is to infer if the given range query contains any training point.
+In this paper, we denote the “neighborhood” as the “range”, and its corresponding distance function as the “range function”.
 Range membership inference attacks extend the formulation of membership inference attacks.
@@
 This extended formulation produces composite hypotheses in the likelihood ratio tests, which are the standard and best attack techniques in MIAs [39, 45, 5, 48].
-Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests and Bayes factors.
+Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests (GLRTs) and Bayes factors.
 We show that RaMIAs can provide a more comprehensive notion of privacy by detecting private information leakage from the vicinity of training data when MIAs underestimate such privacy risk.
-Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 0.20 if we test image classifiers with horizontally flipped images (Figure 2(c)).
-RaMIA, implemented with our simple attack strategy (Sec 4), supersedes MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
+Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 20% if we test image classifiers with horizontally flipped images (Figure 2(c)).
+This is actually the expected behavior of MIAs since the transformed training data are non-members by definition.
+RaMIA, implemented with our simple attack algorithm (Sec 4), surpasses MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
 In this paper, we emphasize the motivation and formulation of our newly proposed attack framework, RaMIA.
-As a proof-of-concept, we experiment RaMIA with a simple attack strategy on tabular, image and text datasets, where RaMIA unanimously outperforms MIA.
+As a proof-of-concept, we implement RaMIA with a simple attack strategy and experiment on tabular, image and text datasets, where RaMIA unanimously outperforms MIA.
 Additionally, our attack can also be potentially used in the pioneering membership inference and data extraction attacks on generative models [4, 44, 7], where the current evaluation requires finding the closest training image for all candidate data and computing their distances.
@@
 (c) MIAScore=1.8 and 1.0
-Figure 1: Examples of similar data with drastically different MIA scores. MIA would classify one of them as member, but the other as non-member.
+Figure 1: Examples of similar data with drastically different MIA scores.
+MIA would classify one of them as member, but the other as non-member.
+The dog image is from the CIFAR-10 dataset, while the text data is from AG News.
+The image classifer is not trained with horizontal flipping augmentation.
 2 Preliminaries
@@
 Shokri et al.
-[39] use a shadow model based approach where shadow models are trained on known training sets in similar ways to the target model.
+[40] use a shadow model-based approach where shadow models are trained on known training sets in similar ways to the target model.
 Confidence values of the training and test data on the shadow models are computed, which are then used as benchmarks in testing.
@@
 To further boost the attack power, Zarifzadeh et al.
-[47] assume the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
-Recent attacks [4, 47] further boost the attack performance on image data by augmenting the test queries with train-time augmentations.
-This assumes that the attacker knows the exact train-test augmentations in advance, and is able to sample from them.
-Augmenting training images with non train-time augmentations is not considered for valid reasons: those augmented images would be non-members in the current privacy notion.
+[48] assumes the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
+Membership inference attacks with augmentations
+Recent attacks [5, 48] find extra attack performance on image data from augmenting the test queries with train-time augmentations, as these augmented training images could have been seen by the model during training.
+However, this assumes that the attacker knows the exact train-time augmentations in advance, and is able to sample from them.
+Augmenting training images with non train-time augmentations is not considered and allowed for valid reasons: those augmented images would be non-members in the current privacy notion based on membership inference, and testing them as members would violate the current privacy notion.
 2.2 Range queries
-If we make a connection to the field of databases, a membership inference attack operates on point queries or exact match queries.
+If we make a connection to the field of databases, the existing membership inference attack operates on point queries or exact match queries.
 That is, each query to the membership inference attack only contains one data point and the attack only concerns if this very point is in the training set.
-On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the "range".
+On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the ”range”.
 The most fundamental difference to point query is that the retrieved result often contains multiple data points instead of a single one.
@@
 By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC).
-A clueless adversary who can only randomly guess the membership labels will get an AUC of 0.5.
-For stronger adversaries, they predict membership more correctly at each error level.
+A clueless adversary who can only randomly guess the membership labels is expected to get an AUC of 0.5.
+For stronger adversaries, they predict membership more accurately at each error level.
 Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
@@
 Figure 2 shows the MIAs under-perform on non-original data.
-This inspires our formulation of RaMIA, where these points will be classified as "members" for better and more comprehensive privacy auditing.
+This inspires our formulation of RaMIA, where these points will be classified as ”members” for better and more comprehensive privacy auditing.
 (a) Purchase-100
@@
 In 2(b), the point query changes to photos of the same identity who has at least one photo in the training set.
-In 2(c), the queries are vertically flipped images.
+In 2(c), the queries are horizontally flipped images.
 In 2(d), $d$ is the Hamming distance to original sentences.
@@
 The main difference between the two games is that the adversary now receives a range query (Step 4 in Def 2) instead of a point query (Step 3 in Def 1).
-We assume that the adversary is able to sample within the range.
-This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [39, 44, 47].
+We assume that the adversary is able to sample a set of points covered by any given range, which effectively means sampling within the range.
+This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [40, 45, 48] as specified in the membership inference game formulation.
 Given a range query and a sampler of $\pi$, it is not difficult to sample within the range.
@@
 One way to visualize a range is to imagine a unit $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions.
-Our framework can cater to any arbitrary range function.
-It can be spatial based (e.g.
+Our framework can cater to any arbitrary range function that is potentially privacy-preserving.
+It can be spatially based (e.g.
 $l_{p}$ distances), transformation based (e.g.
-geometric transformations) and semantic based (e.g.
+geometric transformations), and semantic based (e.g.
 owner/main features of the data).
 In the experiment section, we will present results with all of these types of range functions.
-Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based.
+Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based, which means the privacy concern is whether any other data from the same owner are used in training.
+We will comment on the relations between RaMIA and other existing inference attacks later in the paper.
 How to construct a range
 In Step 3 of the range membership inference game, the details of how the challenger samples the ranges are intentionally omitted.
-This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both in- and out-ranges.
-The details of how we construct the in- and out-ranges for our experiments are elaborated in Appendix B.
+This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both IN and OUT-ranges.
+The details of how we construct the iIN and OUT-ranges for our experiments are elaborated in Section 5.1.
 3.3 Evaluation of RaMIA
-Similarly, we evaluate RaMIA with AUCs.
+Similar to MIAs, we evaluate RaMIA with AUCs.
 However, the notion of true positives and false positives are different from those defined in MIA, as both are defined on the range level.
-To avoid confusion, we call them Range TPR and Range FPR, which means a range is correctly/wrongly predicted to contain at least one training point.
+Specifically, if a range contains at least one training point, it is IN; otherwise, it is OUT.
+The TPR in RaMIA is the proportion of IN-ranges that are correctly identified by the attacker.
+On the other hand, the FPR in RaMIA is the proportion of OUT-ranges that yield positive predictions by the attacker.
+To avoid confusion, we call them (Range) TPR and (Range) FPR.
 4 Range membership inference attacks
+4.1 (Simple) Hypothesis testing
+The standard way to tackle the inference game (Def 1) is to apply statistical hypothesis tests [45, 5]:
+$H_{0}$ $:\text{The given }z\text{ is not a training point }(b=0).$
+$H_{1}$ $:\text{The given }z\text{ is a training point }(b=1).$
+The likelihood ratio test (LRT) is then conducted
+$\frac{\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.49994ptP}{\left(\theta|H_{0}\right)}}$ (1)
+This is usually called ”simple” hypothesis testing because each $H$ contains a single hypothesis.
+The scoring function in the membership inference attack can be considered as an approximation of the likelihood function $\text{I\kern-1.49994ptP}{(\cdot)}$.
 4.2 Composite hypothesis testing
-Similar to likelihood ratio tests for membership inference games A.1, we can also construct two hypotheses for range membership inference game (Def 2):
+Similar to likelihood ratio tests for membership inference games 4.1, we can also construct two hypotheses for range membership inference game (Def 2):
 […]
-$\forall z\in\mathcal{R}_{b}:z\not\in D.$
 […]
-$\exists z\in\mathcal{R}_{b}\text{ s.t. }z\in D.$
 […]
-To make full use of the Bayes Factor, we need to know the prior distribution, which is unrealistic.
-On the other hand, taking the max seems to be a more intuitive approach for range membership inference attacks, because it provides a two-step solution: search and test.
-Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points.
-However, this assumes that we can reliably find the max values in a given range.
+[…]
+Since it can be intractable to iterate over all points in a given range $\mathcal{R}_{b}$, we use a set of sampled points $S$ in the given range as its proxy. In this way, the two hypotheses are:
+[…]
+[…]
+[…]
+[…]
+The likelihood ratio in this case is still $\frac{\text{I\kern-1.04996ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.04996ptP}{\left(\theta|H_{0}\right)}}$.
+However, the alternative hypothesis $H_{1}$ is composite because it is a union of multiple hypotheses $\bigcup_{z_{i}\leftarrow S}(z_{i}\in D)$.
+Therefore, we need to use statistical methods tailored for composite hypothesis testing.
+There are two commonly used methods for it: Bayes Factor [22] and Generalized Likelihood Ratio Tests (GLRTs) [43].
+GLRT
+GLRT assumes that the true hypothesis $h^{*}$ is explicitly listed in the composite hypothesis, which, in our case, means the training point is in the set of samples $\{z_{i}\}$.
+In this case, the composite hypothesis can be reduced to a single hypothesis by taking the maximum:
+$\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}\Rightarrow\max_{x\in S}\text{I\kern-1.49994ptP}{(\theta|x\in D)}.$ (2)
+Bayes Factor
+Bayes Factor assumes the hypothesis is a random variable, each with a likelihood value according to a prior distribution.
+In our case, it means each point $z$ in the range can be sampled with probability determined by the prior data distribution $\text{I\kern-1.49994ptP}{(z)}$.
+To reduce the composite hypothesis to a simple one, Bayes Factor computes the weighted average:
+[…]
+This can be thought of as taking the expectation of the composite hypothesis or using the most representative hypothesis instead of the maximal (as in GLRT) in the likelihood ratio test.
+Why both methods fall short in RaMIA
+At first look, GLRT might be an intuitive choice, because taking the maximum leads to a two-step strategy: search and test.
+Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points, and their membership should be indicative of the ranges’ membership.
+However, this assumes the true training point is included in every sampled set created by the adversary, which is extremely unlikely.
+Secondly, this also assumes that we can reliably find the max values in a given range.
 Since most ranges are large data subspaces, it is very challenging to find the optimal points within the large space.
 Even if the search space can be navigated, any search algorithm is likely to return local maxima.
-Hence, a robust way is to aggregate the top samples.
-However, membership inference attacks are known to be unreliable on out-of-distribution (OOD) data [47] and can assign them high scores.
+To make full use of the Bayes Factor, we need to know the prior distribution, which is also not possible.
+A more important issue is that the power of existing statistical methods relies on the absolute correctness of the likelihood values: the larger the likelihood values, the more likely the point is a member.
+In other words, the largest likelihood values of non-members must be smaller than the smallest likelihood of members.
+However, membership inference attacks are known to be non-perfect; they are also known to be particularly unreliable on out-of-distribution (OOD) data [48] and can assign them high scores.
 When the sampling space contains only these data as opposed to real and in-distribution (ID) data, the maximum might not be anything close to true training data, increasing the FPR and lowering the AUC as a result.
-In this paper, we adopt a simple attack strategy.
-Our solution to this is based on the type of data in the sampling space.
-If the data are all naturally ID, we can combine Bayes Factor and GLRT by taking the average likelihood of the top samples to reduce the influence of the randomness in the sampling process.
-On the other hand, if the adversary can only synthesize data within the range, the top samples are highly likely to be OOD data with high membership scores.
+4.3 Our approach: Trimmed averages
+Given these limitations, in this paper, we propose a simple attack strategy based on Bayes Factor.
+We first assume every sampled point is of equal probability.
+Then, to address the unreliability of MIA scores, we use a trimming process based on the type of data in the sampling space.
+If the data are all naturally ID, we trim the less probable samples and take the average likelihood of the top samples to reduce the influence of non-members and the randomness in the sampling process.
+On the other hand, if the adversary can only synthesize data within the range instead of sampling from a pool of real data, the top samples are highly likely to be OOD data with high membership scores.
 Hence, in this case, we want to remove those points from the equation.
-Since the presence of training points intuitively raises the average membership score of ID points nearby, compared to having no training points at all, we average the membership scores of the remaining samples.
-Unifying these two strategies gives us the following:
-$\mathbb{P}\left(\theta|H_{1}\right)=\text{TrimmedAvg}(S,q_{s},q_{e};\mathbb{P})=\text{Avg}_{x\not\in[q_{s},q_{e}]\text{-th quantiles}}\mathbb{P}(\theta|x\in D),$ (1)
-where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles where we want to remove to compute our robust statistics that are one-sided trimmed means.
+We then trim the top-scoring samples and take the average of the rest.
+This gives us the following:
+$\end{multlined}\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}=\text{TrimmedAvg}(S,q_{s},q_{e};\text{I\kern-1.49994ptP})\\ =\text{Avg}_{x\not\in[q_{s},q_{e}]\text{-th quantiles}}\text{I\kern-1.49994ptP}{(\theta|x\in D)},$ (4)
+where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles that we want to remove to compute our robust statistics that are one-sided trimmed means.
 If the sampling space is filled with synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points.
@@
 $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available.
+If we do not trim anything, the formulation reduces to Bayes Factor which assigns equal probability to all points.
 Note that the optimal hyperparameters may differ across different membership signals (e.g.
 loss values, LiRA scores), as they exploit different vulnerabilities and expose different training points.
-However, for fixed model architectures, range functions, data distributions and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [47].
-Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters.
+However, for fixed model architectures, range functions, data distributions, and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [48].
+Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters in a grid search manner.
+More details are in Appendix A.
 4.4 Range membership inference attack as a framework
 The range membership inference attack is a new inference attack framework, not a particular attack algorithm.
-There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 1.
+There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 4.
 The sampler $\text{Sampler}(\mathcal{R}):S\rightarrow X$ returns samples within the given range.
-The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\mathbb{P}(\theta|x)$.
-A number of existing attack algorithms can be plugged in.
+The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\text{I\kern-1.49994ptP}{(\theta|x)}$.
+Any existing MIA algorithm can be plugged in as RaMIA’s membership testing backbone.
 Similar to MIAs, the key to using RaMIA as a privacy auditing tool is to compute the range membership score.
@@
 We experiment on the commonly used Purchase-100 [40], CelebA [28], CIFAR-10 [25] and AG News [49] datasets.
-Details of how we split the dataset, train models, construct ranges, and obtain samples are explained in Appendix B.
-We compare RaMIA with MIA (Sec 5.1) in scenarios MIA under-performs (Depicted in Figure 2).
-Both attacks are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [47], with three reference models trained in the same way as Carlini et al.
-[5] and Zarifzadeh et al.
-[47].
+We will describe the details of how we split the dataset, train models, construct ranges, and obtain samples in Section 5.1.
+Since the range membership notion is new, we do not have a prior method to compare with.
+However, since the goal of introducing this new privacy notion is to enable better and more comprehensive privacy auditing, we compare RaMIA with the de-factor privacy auditing framework, MIA.
+We show that RaMIA performs better (Sec 5.5) in scenarios where MIA under-performs (Depicted in Figure 2).
+Both MIA and RaMIA in the experiments are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [48], with a few reference models trained in the same way as [5] and [48].
 The respective queries are outlined in Table 1, and the definitions of members under each attack framework are explained in Table 2.
@@
 AG News sentences that are of Hamming distance 8 to sentence $x^{\prime}$ sentence $x^{\prime}$
-Table 2: Definitions of range and point members corresponding to the attack queries in Table 1
+Table 2: Definitions of range and point members corresponding to the attack queries in Table 1.
 Dataset Range member if there is at least (Point) member if
@@
 AG News one training sentence within Hamming distance 8 to $x^{\prime}$ $x^{\prime}$ is member
-On Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$.
+5.1 Setup details
+As mentioned earlier, the range function must come from the semantics of privacy.
+Hence, in experiments, we use specific range functions with different data types.
+For tabular data, we consider missing columns, which is an extreme case of using Euclidean distance on missing columns as a range function.
+For human photo data, we use a semantic range function, which is based on the main person featured in the photo.
+For other image data, we use geometric transformations as range functions.
+For text data, we use (word level) Hamming distance, which is edit distance or Levenshtein distance that only considers word substitution.
+The reasons for choosing these range functions have been motivated and explained in earlier parts of this paper.
+5.1.1 Tabular data: Purchase-100
+  • Dataset Purchase-100 [40] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data.
+This dataset was first curated by [40] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product.
+The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
+  • Models We train five-layer multi-layer perceptron (MLP) models in PyTorch [34] on half of the entire dataset.
+The hidden layers are of sizes $[512,256,128,64]$, identical to the setup in [48].
+All models achieve a test accuracy of $83\%$.
+  • Construction of ranges We simulate the scenario where the attacker has incomplete data (data with missing values).
+For all training and test data records, we randomly mask k columns.
+Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary.
+We then check if each range constructed from test data points includes any training point.
+If so, they are re-labeled as ”IN-ranges”.
+  • Sampling within ranges Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns.
+The parameter of the sampler is computed by taking the average value of each column.
+Because of the nature of this dataset, our sampled data can be regarded as in-distribution.
+We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
+5.1.2 Image data I: CelebA
+  • Dataset CelebA [28], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features.
+We construct the members set by only including photos of celebrities with identity numbers smaller than 5090.
+The rest are used to construct the non-members set.
+For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
+  • Models We train four-layer convolutional neural networks (CNNs) in PyTorch [34] on the training set to predict the facial attributes of any given photo.
+Our target model has a test accuracy of $87\%$.
+  • Construction of ranges The range function here is a semantic one that is based on the identity of the face image.
+For example, a range query can be ”all Alice’s photos”.
+Since the identities in the training and non-members set are disjoint, it is easy to construct IN- and OUT-ranges based on the distribution of identities in the two sets.
+  • Sampling within ranges For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
+5.1.3 Image data II: CIFAR-10
+  • Dataset CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
+  • Models We train WideResNets-28-2 [47] with JAX [2] on half of the training set of CIFAR-10 using the code from [5], with and without image augmentations.
+Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$.
+The train time augmentation is the composition of random flipping, cropping and random hue.
+  • Construction of ranges The range function here is different types of image augmentations, which are geometric transformations.
+An example of a range query is ”all transformed version of image $X$”.
+For each training and test image, a range is constructed by applying different transformations.
+  • Sampling within ranges For each range query, we independently apply 15 image augmentations on the query center.
+The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
+(a) Purchase-100 (m=10)
+(b) CelebA
+(c) CIFAR-10
+(d) AG News (d=5)
+Figure 3: RaMIAs perform better than MIAs on points that are close to original points but not exactly the same.
+5.1.4 Textual data: AG News
+  • Dataset We use the AG News dataset [49], which is a news collection with four categories of news.
+Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset.
+There are 120,000 sentences in its training set.
+It is a popular text dataset in LLM MIA literature [33].
+  • Models We took pretrained GPT-2 [35] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library.
+The finetuning is done for 4 epochs.
+Our target model achieves a perplexity of 1.39 on the test set of AG News.
+  • Construction of ranges The range function here is word-level Hamming distance, which can be thought of as the edit distance measured on word level that only allows word substitution.
+An example of a range query is ”all sentences within Hamming distance $d$ to sentence $x$”.
+To construct IN- and OUT-ranges, we just need to specify the max Hamming distance and the starting sentence.
+We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pretrained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences.
+A Hamming distance is then specified with each starting sentence to form a range.
+  • Sampling within ranges We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
+5.2 Hyperparameters
+Overall, on Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$.
 On CIFAR-10, we apply up to 15 distinct transforms, and set $q_{e}=100,q_{s}=40$.
@@
 We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
+5.3 Implementation details
+We train 8 models on Purchase-100 and CIFAR-10, and train 4 models on CelebA and AG News.
+Each model is trained on half of the dataset, as described in [5, 48].
+For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001.
+For WideResnets, we use the training code from [5].
+On AG News, the models are trained for 4 epochs.
+On other datasets, they are trained for 100 epochs.
+All training is done on two Nvidia RTX 3090 GPUs.
+Training on AG News takes about 1 hour per epoch.
+Training other models takes less than one hour each.
+5.4 Metrics
+To evaluate the performance of RaMIA and MIA based on the same MIA algorithm, we use AUCs.
+The inputs to RaMIA are range queries in Table 1, while the inputs to MIA are point queries, which are range centers, in Table 1.
+The ground truth membership for both RaMIA and MIA is the range membership information defined in Table 2, which allows us to compare the power of both inference attack frameworks in identifying queries that are privacy leaking.
+In other words, both point and range TPRs/FPRs are calculated against the range membership information.
+Following [5], we also report (Range) TPR at small (Range) FPR for both MIA and RaMIA (Table 4).
 5.5 RaMIAs quantify privacy risks more comprehensively than MIAs
@@
 50 sentences are too little to meaningfully cover anything in the space.
-But yet limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is flawed and needs a better framework.
-Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 4.TPRs at small FPRs are in Table 3
-(a) Purchase-100 (m=10)
-(b) CelebA
-(c) CIFAR-10
-(d) AG News (d=5)
-Figure 3: RaMIAs perform better than MIAs on points that are close to original points but not exactly the same.
+Yet, limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is suboptimal and needs a better framework.
+Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 3.TPRs at small FPRs are in Table 4.
+The improvement in AUC is summarized in Table 3.
+Table 3: Improvement in AUCs after switching from MIA to RaMIA across multiple iterations of random sampling. We do not randomly sample but use all available attack images in CelebA.
+Purchase-100 CIFAR-10 CelebA AG News
+$\Delta\text{AUC}$ $2.62\pm 0.04$ $4.12\pm 0.06$ $5.4$ $1.20\pm 0.2$
+Relation to user-level inference
+Note that in the CelebA experiment, we use identify information as the semantic range function, making it similar to user-level inference.
+This further shows that RaMIA is a better and more comprehensive privacy auditing framework.
+In terms of algorithms, our attack strategy should dominate the simple averaging approach used in prior work [31, 23], since the trimming ratio is optimized (compared to no trimming in simple averaging).
+There are other user-level inference algorithms [11] that train additional shadow models and classifiers, which incurs additional computational costs.
+They also consider scenarios where user information is the label of their target models, e.g.
+facial and speech recognition systems that predict user ID as their outputs [11, 10].
+This encourages their target models to explicitly cluster data based on user information, making inference easier.
+On the other hand, our target model, a facial attribute classifier, does not use user information anywhere in the training, making the inference harder.
+Table 4: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$.
+MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them.
+Standard deviation over random sampling iterations is reported, except for CelebA, where we use all available candidates.
+The TPR and FPR are calculated based on the range membership information, as described in Sec 5.4.
+For a fair comparison, we should compare RaMIA and MIA based on the same membership testing backbone, e.g.
+MIA with RMIA versus RaMIA with RMIA.
+Purchase-100 CIFAR-10 CelebA AG News
+TPR@FPR(%) 1% 0.1% 1% 0.1% 1% 0.1% 1% 0.1%
+MIA
+LOSS 0 0 0.15 0 1.86 0.31 $80.0$
+RMIA 2.18 0.37 2.40 0.21 1.69 0.19 $0.30$ 00
+RaMIA
+LOSS $0\pm 0$ $0\pm 0$ $1.17\pm 0.06$ $0.05\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
+RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $3.59\pm 0.11$ $0.80\pm 0.02$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
 5.6 Factors affecting RaMIA performance
@@
 (Sanity Check)
-(b) RaMIA performs better on CIFAR-10 when the range size increases. The size is equal to the number of distinct transformation applied to images.
+(b) RaMIA performs better on CIFAR-10 when the range size increases. The range size here is equal to the number of distinct transformations applied to images.
 Figure 4: Attack performance increases in tandem with training point density.
+(a) Purchase-100
+(b) CIFAR-10
+(c) AG News
+Figure 5: Correlation between the percentile of RaMIA and MIA scores of members among non-members.
+The larger the percentile is, the more non-members the member dominates, and the more likely for it to be classified as IN.
+The Pearson correlation coefficients are provided for each plot.
+It shows that the vulnerability to MIA is positively correlated with that to RaMIA.
 Susceptibility to MIAs and RaMIAs is correlated
@@
 Researchers have previously discovered that machine learning models memorize duplicate data more [26, 6].
-In our CelebA dataset, each celebrity has different numbers of photos in the training set, which can be thought as that each identity has different levels of duplication in the training set.
+In our CelebA dataset, each celebrity has a different number of photos in the training set, which can be thought of as each identity having different levels of duplication in the training set.
 Similar to the insights from MIAs, we also observe that identities that have more training images, i.e.
 higher duplication rate, are more susceptible to RaMIA.
-Figure 8 shows the relationship between the percentile each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate.
+Figure 6 shows the relationship between the percentile of each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate.
 Generally speaking, identities that have more training photos are more prone to RaMIAs.
-Similarly correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 7.
+Similarly, correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 5.
+Figure 6: Number of photos from the same celebrity in the training set affects the identifiability of the range.
 5.7 Mismatched training and attack data hurts attack performance
@@
 Since they adopt the existing privacy notion based on point queries, only (augmented) images seen by the model in the training stage are considered as members.
-Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy of notion.
+Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy notion.
 In RaMIA, the set of augmentations is given by the challenger (Def 2), which can contain augmentations not used in training, but considered as privacy leaking.
 Using the aggregation method in [5] will hurt the attack performance if non-training augmentations are used.
-However, RaMIA is designed to be robust in this scenario (Fig 3(c)).
+However, RaMIA is designed to be robust in this scenario (Fig 3(c)) due to its trimming process.
+5.8 RaMIA on redacted data
+Many large language models (LLMs) are trained with sensitive textual data.
+Some of the data with sensitive information redacted might be publicly available.
+Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM.
+Accurately identifying the redacted sentences paves the way for reconstructing them as a follow-up attack.
+Figure 7 shows the results.
+In this experiment, we use spaCy [18] to mask peoples’ names to simulate the masking of personally identifiable information (PII).
+We then generate 10 possible sentences for each masked sentence using BERT and conduct RaMIA.
+The MIA performance is the average attack performance over all 10 possible sentences.
+The performance gap is smaller compared to that in Figure 3(d).
+The reason might be that BERT fails to produce diverse PII completions, making all candidate sentences similar to each other, and reducing the power of RaMIA.
+Figure 7: RaMIA on a subset of AG News where names are redacted.
 6 Conclusion
@@
 Moving the query to similar points causes a drastic decrease in performance, rendering MIAs less useful.
-We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs.
-and cover the failure cases of MIAs by checking if a given range contains a training point.
+We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs and covers the failure cases of MIAs by checking if a given range contains a training point.
 We introduce RaMIA as an attack framework that can be implemented with any existing MIA algorithm.
@@
 In future work, we hope to design more powerful RaMIA strategies that are robust to the change of membership signals and datasets, especially on LLMs where we believe our privacy notion is extremely relevant.
-Appendix A Attack Details
-A.1 (Simple) Hypothesis testing
-The standard way to tackle the inference game (Def 1) is to apply statistical hypothesis tests [44, 5]:
-$H_{0}$ $:\text{The given }z\text{ is not a training point }(b=0).$
-$H_{1}$ $:\text{The given }z\text{ is a training point }(b=1).$
-The likelihood ratio test (LRT) is then conducted
-$\frac{\mathbb{P}\left(\theta|H_{1}\right)}{\mathbb{P}\left(\theta|H_{0}\right)}$ (2)
-This is usually called "simple" hypothesis testing because each $H$ contains a single hypothesis.
-A.2 Attack algorithms
+Appendix A Attack algorithms
 In this section, we explain the details of the membership inference attack algorithms used in our experiments.
@@
 LOSS [46] computes loss values as a proxy of membership score on given points: $\text{MIA}(x;\theta)=\mathit{l}(x;\theta)$.
-To compute the likelihood, an easy way is to take the exponential of the negative of the loss $P=\exp^{-\mathit{l}}$.
+To compute the likelihood, an easy way is to take the exponential of the negative of the loss $\text{I\kern-1.49994ptP}=\exp^{-\mathit{l}}$.
 RMIA
-RMIA [47] computes membership score by applying chain rule: $\mathbb{P}(\theta|x)=\frac{\mathbb{P}(x|\theta)\mathbb{P}(\theta)}{\mathbb{P}(x)}$.
-The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\mathbb{P}_{z\in Z}[\frac{\mathbb{P}(\theta|x)}{\mathbb{P}(\theta|z)}\leq\gamma]$, where the term $\mathbb{P}(\theta)$ will cancel out with each other.
-The normalizing constant $\mathbb{P}(x)$ is computed with reference models: $\mathbb{P}(x)=0.5\mathbb{E}_{\theta_{\text{in}}}\mathbb{P}(x|\theta_{\text{in}})+0.5\mathbb{E}_{\theta_{\text{out}}}\mathbb{P}(x|\theta_{\text{out}})$.
+RMIA [48] computes membership score by applying chain rule: $\text{I\kern-1.49994ptP}{(\theta|x)}=\frac{\text{I\kern-1.04996ptP}{(x|\theta)}\text{I\kern-1.04996ptP}{(\theta)}}{\text{I\kern-1.04996ptP}{(x)}}$.
+The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\text{I\kern-1.49994ptP}_{z\in Z}{[\frac{\text{I\kern-1.04996ptP}{(\theta|x)}}{\text{I\kern-1.04996ptP}{(\theta|z)}}\geq\gamma]}$, where the term $\text{I\kern-1.49994ptP}{(\theta)}$ will cancel out with each other.
+The normalizing constant $\text{I\kern-1.49994ptP}{P(x)}$ is computed with reference models: $\text{I\kern-1.49994ptP}{(x)}=0.5\mathbb{E}_{\theta_{\text{IN}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{IN}})}+0.5\mathbb{E}_{\theta_{\text{OUT}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{OUT}})}$.
 In its offline version, the in models are unavailable.
-In this case, the former probabilities is approximated by the latter term $\mathbb{P}_{\text{in}}=a\mathbb{P}_{\text{out}}+(1-a)$.
+In this case, the former probabilities are approximated by the latter term $\text{I\kern-1.49994ptP}_{\text{IN}}=a\text{I\kern-1.49994ptP}_{\text{OUT}}+(1-a)$.
 The hyperparameter $\alpha$ is chosen based on the reference models.
 Specifically, one reference model is chosen as the temporary target model, and the rest are used to attack it.
-The value of $\alpha$ is chosen to be the best performing value under this setting, obtained via a simple sweeping.
+The value of $\alpha$ is chosen to be the best-performing value under this setting, obtained via a simple sweeping.
 In our experiment, we use the offline attack only.
@@
 For AG News, we set it to be 1.0.
-Appendix B Setup Details
-On each of the dataset, we train four models on half of the dataset in the same way described by Carlini et al.
-[5], Zarifzadeh et al.
-[47].
-We will describe the details of the datasets below.
-We have checked with their licenses with our best effort, and confirm their terms of use are respected.
-B.1 Tabular data: Purchase-100
-Dataset
-Purchase-100 [39] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data.
-This dataset was first curated by Shokri et al.
-[39] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product.
-The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
-Models
-We train five-layer multi-layer perceptron (MLP) models in PyTorch [33] on half of the entire dataset. The hidden layers are of sizes $[512,256,128,64]$. All models achieve a test accuracy of $83\%$.
-Construction of ranges
-We simulate the scenario where the attacker has incomplete data (data with missing values).
-For all training and test data records, we randomly mask k columns.
-Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary.
-We then check if each range constructed from test data points include any training point.
-If so, they are re-labelled as "in-ranges".
-Sampling within ranges
-Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns.
-The parameter of the sampler is computed by taking the average value of each column.
-Because of the nature of this dataset, our sampled data can be regarded as in-distribution.
-We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
-B.2 Image data I: CelebA
-Dataset
-CelebA [28], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features.
-We construct the members set by only including photos of celebrities with identity number smaller than 5090.
-The rest are used to construct the non-members set.
-For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
-Model
-We train four-layer convolutional neural networks (CNNs) in PyTorch [33] on the training set to predict the facial attributes of any given photo. Our target model has a test accuracy of $87\%$.
-Construction of ranges
-The range function here is a semantic one that is based on the identity of the face image.
-For example, a range query can be "all Alice’s photos".
-Since the identities in the training and non-members set are disjoint, it is easy to construct in- and out-ranges based on the distribution of identities in the two sets.
-Sampling within ranges
-For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
-B.3 Image data II: CIFAR-10
-Dataset
-CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
-Models
-We train WideResNets-28-2 [46] with JAX [2] on half of the training set of CIFAR-10 using the code from [5], with and without image augmentations.
-Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$.
-The train time augmentation is the composition of random flipping, cropping and random hue.
-Construction of ranges
-The range function here is different types of image augmentations, which are geometric transformations.
-An example of a range query is "all transformed version of image $X$".
-For each training and test image, a range is constructed by applying different transformations.
-Sampling within ranges
-For each range query, we independently apply 15 image augmentations on the query center.
-The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
-B.4 Textual data: AG News
-Dataset
-We use AG News dataset [48], which is a news collection with four categories of news.
-Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset.
-There are 120,000 sentences in its training set.
-Model
-We took pretrained GPT-2 [34] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library.
-The finetuning is done for 4 epochs.
-Our target model achieves a perplexity of 1.39 on the test set of AG News.
-Construction of ranges
-The range function here is word-level Hamming distance, which can be thought as the edit distance measured on word level that only allows word substitution.
-An example of a range query is "all sentences within Hamming distance $d$ to sentence $x$".
-To construct in- and out-ranges, we just need to specify the max Hamming distance and the starting sentence.
-We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pre-trained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences.
-A Hamming distance is then specified with each starting sentence to form a range.
-Sampling within ranges
-We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
-Appendix C Implementation details
-For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001.
-For WideResnets, we use the training code from [4].
-On AG News, the models are trained for 4 epochs.
-On other datasets, they are trained for 100 epochs.
-All training are done on two Nvidia RTX 3090 GPUs. Training on AG News takes about 1 hour per epoch. Training other models takes less than one hour each.
-Appendix D RaMIA on redacted data
-Many large language models (LLMs) are trained with sensitive textual data.
-Some of the data with sensitive information redacted might be public available.
-Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM.
-Accurately identify the redacted sentences paves the way for reconstructing them as a follow-up attack.
-Figure 5 shows the results.
-In this experiment, we use spaCy [18] to mask peoples’ names to simulate masking of personally identifiable information (PII).
-We then generate 10 possible sentences for each masked sentence using BERT, and conduct RaMIA.
-The MIA performance is the average attack performance over all 10 possible sentences.
-Figure 5: RaMIA on subset of AG News where names are redacted.
-Appendix E Extra results
-In this section, we put extra experiment results.
-Table 3: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$.
-MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them.
-Standard deviation over random sampling iterations are reported, except for CelebA, where we use all available candidates.
-Purchase-100 CIFAR-10 CelebA AG News
-TPR@FPR(%) 1% 0.1% 1% 0.1% 1% 0.1% 1% 0.1%
-MIA
-LOSS 0 0 0.15 0 1.86 0.31 $80.0$
-RMIA 2.18 0.37 2.40 0.21 1.69 0.19 $0.30$ 00
-RaMIA
-LOSS $0\pm 0$ $0\pm 0$ $1.17\pm 0.06$ $0.05\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
-RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $3.59\pm 0.11$ $0.80\pm 0.02$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
-Table 4: Improvement in AUCs after switching from MIA to RaMIA across multiplue iterations of random sampling. We do not random sample but use all available attack images in CelebA.
-Purchase-100 CIFAR-10 CelebA AG News
-$\Delta\text{AUC}$ $2.62\pm 0.04$ $4.12\pm 0.06$ $5.4$ $1.20\pm 0.2$
-Figure 6: RaMIA gets better when the range size increases on CelebA, even if we assume the attacker cannot sample any training image.
-(a) Purchase-100
-(b) CIFAR-10
-(c) AG News
-Figure 7: Correlation between the percentile of RaMIA and MIA scores of members among non-members.
-The larger the percentile is, the more non-members the member dominates, and the more likely for it to be classified as IN.
-The Pearson correlation coefficients are provided for each plot.
-It shows that the vulnerability to MIA is correlated with that to RaMIA.
-Figure 8: Number of photos from the same celebrity in the training set affects the identifiability of the range.

### v3

@@
 Abstract
-Machine learning models can leak private information about their training data, but the standard methods to measure this risk, based on membership inference attacks (MIAs), have a major limitation.
-They only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping data revealing the same private information.
+Machine learning models can leak private information about their training data.
+The standard methods to measure this privacy risk, based on membership inference attacks (MIAs), only check if a given data point exactly matches a training point, neglecting the potential of similar or partially overlapping memorized data revealing the same private information.
 To address this issue, we introduce the class of range membership inference attacks (RaMIAs), testing if the model was trained on any data in a specified range (defined based on the semantics of privacy).
@@
 We show that RaMIAs can capture privacy loss more accurately and comprehensively than MIAs on various types of data, such as tabular, image, and language.
-RaMIA paves the way for a more comprehensive and meaningful privacy auditing of machine learning algorithms.
-1 Introduction
+RaMIA paves the way for more comprehensive and meaningful privacy auditing of machine learning algorithms.
+I Introduction
 Machine learning models are prone to training data memorization [1, 2, 3, 4, 5].
 It is also a known fact that the outstanding predictive performance of machine learning models on long-tailed data distributions often comes at the expense of blatant memorization of certain data points [2, 6, 7, 8].
-In simple words, memorization is the phenomenon that models behave differently on training points, compared to other points.
-The memorization can lead to significant privacy risks as adversaries can infer private information about the training data from only having black box access to models.
-To quantify the privacy risk of machine learning models, a privacy notion needs to be agreed upon first.
-The reigning privacy notion is defined by membership information.
-Membership information of a data point is binary, but this single bit of information carries huge privacy implications.
-Being able to infer membership information opens up the possibility of conducting data reconstruction attack [37, 17, 4, 29], where the reconstruction attack inspects the membership of plausible data points to recover the training set.
-The de-facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [40], where an adversary aims to predict whether a given query data is part of the training set of the target model.
+Memorization refers to the phenomenon where models behave differently on data depending on whether it was included in the training set.
+Such behavior can lead to significant privacy risks because adversaries can infer sensitive information about training data even with only black-box access to the model.
+To quantify the privacy risk of machine learning models, it is crucial to define a precise privacy notion.
+The prevailing privacy notion is based on membership information, a binary indicator that carries substantial privacy implications.
+Accurate inference of membership status can enable data reconstruction attacks [9, 10, 11, 12], where the adversary probes the membership of plausible data points to recover the training set.
+The de facto way to audit the privacy risk according to this privacy notion is to conduct membership inference attacks (MIAs) [13], where an adversary aims to predict whether a given point belongs to the training set of the target model.
 The more powerful the membership inference attack is, the higher the privacy risk the target model bears.
-Membership inference attacks provide a lower bound of the true privacy risk of a model, so improving the attack performance also means tightening the bound of privacy risk estimation.
-So far, the community has put much effort into improving the power of membership inference attacks by crafting better membership signals and constructing better statistical tests [39, 40, 36, 45, 5, 48].
-While these have been useful for the betterment of privacy auditing, they have ignored the fundamental drawback of membership inference attacks as a practical privacy auditing tool, i.e., MIAs assume it is a privacy concern only if the adversary can identify the exact, full version of the training data.
-However, if the adversary can identify data points that are similar enough to the training points, it should also be treated as a significant privacy risk, because those points can contain similar levels of private information.
-For example, two Alice’s photos taken from slightly different angles, or with a different background would contain similar private information about Alice’s face or location.
-Similar to geometric transformations in images, small perturbations or rephrasing in textual data also affect little in the sensitivity of the information conveyed [13].
-This oversight means private information leakage beyond exact matches of training data is currently ignored and the contemporary privacy auditing tools might produce overly optimistic results.
-Besides, focusing on exact membership inference attacks renders them incapable of handling queries with missing values.
-This is another major limitation of MIAs, as data records with the same sensitive features and a few missing non-private features could carry a similar level of private information as the full data records.
-Imagine the case where an adversary can infer that an Asian person of age 25, identification number 123456 is in an HIV patients dataset, there is no need to identify or collect the exact values of the rest of the features, because the adversary is already able to pinpoint who has HIV from the given subset of key features.
-Even if the key identifiers are missing and unknown, it is still a grave privacy threat if the attacker can infer the rest of the features from the subset of known features that often contain quasi-identifiers, which has been extensively studied in the failure of k-anonymity as a privacy-preserving measure [38], where the attacker can reconstruct and identify training data with certain features masked.
-With existing inference attacks, there are two types of approaches to work with incomplete data.
-The first approach is to impute the missing features before performing membership inference attacks.
-However, if we simply make up the missing values or pass noisy data to the membership inference attack, the attack is expected to output “not a member,” since the chance of us coming up with the right values is extremely slim.
-If we use attribute inference [20] to help fill in the missing features, the imputed data may have inflated membership scores even if they are non-members, leading to a higher false positive rate.
-Therefore, existing MIA-based privacy auditing frameworks does not provide a natural solution to quantifying privacy leakage with missing features.
-The second approach that works with incomplete data is property inference [41, 1, 8, 9, 46], but it aims to extract statistical information about the features instead of testing membership information of data records.
-Hence, none of the existing inference attack frameworks can reliably quantify privacy leakage with incomplete data entries.
-We argue that privacy quantification should not be point-based, because a small neighborhood around training points, defined based on a distance function out of the semantics of privacy, also contains similar private information.
-Hence, in this paper, we are proposing a new inference attack framework called range membership inference attacks (RaMIAs) to better capture the notion of privacy.
-Instead of using point queries and testing for exact matches, range membership inference attacks use range queries that cover a set of points.
-The goal of range membership inference attacks is to infer if the given range query contains any training point.
-In this paper, we denote the “neighborhood” as the “range”, and its corresponding distance function as the “range function”.
-Range membership inference attacks extend the formulation of membership inference attacks.
-We adapt the original inference game formulation to reflect the change to range queries in RaMIAs.
-This extended formulation produces composite hypotheses in the likelihood ratio tests, which are the standard and best attack techniques in MIAs [39, 45, 5, 48].
-Our method is based on standard statistical methods for composite hypothesis testing, namely generalized likelihood ratio tests (GLRTs) and Bayes factors.
-We show that RaMIAs can provide a more comprehensive notion of privacy by detecting private information leakage from the vicinity of training data when MIAs underestimate such privacy risk.
-Specifically, we observe a simple flipping can cause the membership score to decrease from very high to 0 (Figure 1), and the overall AUC can drop 20% if we test image classifiers with horizontally flipped images (Figure 2(c)).
-This is actually the expected behavior of MIAs since the transformed training data are non-members by definition.
-RaMIA, implemented with our simple attack algorithm (Sec 4), surpasses MIA by at least 5% on image datasets (Fig 3(b), 3(c)), providing better privacy auditing at the cost of as few as 15 samples, which is insignificant compared to the dimensionality of the data space.
-In this paper, we emphasize the motivation and formulation of our newly proposed attack framework, RaMIA.
-As a proof-of-concept, we implement RaMIA with a simple attack strategy and experiment on tabular, image and text datasets, where RaMIA unanimously outperforms MIA.
-Additionally, our attack can also be potentially used in the pioneering membership inference and data extraction attacks on generative models [4, 44, 7], where the current evaluation requires finding the closest training image for all candidate data and computing their distances.
-By setting the distance function as the range function and conducting RaMIAs, we can evaluate the attacks more systematically.
+Figure 1: Illustration of privacy notions for MIAs and RaMIAs.
+LEFT: Sample CIFAR-10 training and test data are depicted with one highlighted IN range and one OUT range for visualization.
+In MIAs, only exact training points are deemed private, while all other points in the data space are considered non-private.
+The privacy notion is hence akin to a Dirac delta function that assigns privacy only to exact training points and zero everywhere else.
+RIGHT: RaMIA expands this definition by treating points that are similar to training data as privacy leaking.
+This approach captures privacy risks not only at exact training points but also in their vicinity, offering a more comprehensive and realistic privacy assessment.
+Membership inference attacks provide a lower bound of the model’s true privacy risk, so improving the attack performance also tightens the bound of privacy risk estimation.
+So far, to more accurately audit the privacy risk, the community has largely focused on enhancing MIAs by developing stronger membership signals and more sophisticated statistical tests [14, 13, 15, 16, 17, 18].
+While these advances have improved privacy auditing, they all quantify privacy risks by testing memorization of the exact, full version of training points.
+However, information leakage is not an all or nothing phenomenon.
+In many cases, a model may not memorize a training record in its entirety but rather partial information in the form of key features or data patterns.
+These memorization behaviors, which are more realistic and prevalent, lead to a form of leakage of sensitive information that can be exhibited from similar points in the vicinity of training data.
+For example, a model trained on images including people’s faces may capture the distinctive facial attributes that uniquely identify an individual while ignoring extraneous details such as the background.
+This could be a concern, as an image of the same person taken from a different angle or with a different background has a significant overlap in private information.
+Membership inference attacks, designed to detect exact matches, are not equipped to quantify this notion of information leakage as any slight change to training points would turn them into non-members where any correct MIA is expected to produce negative outcomes.
+For example, when using MIA, a simple horizontal flip can reduce a membership score from a high value to zero (Figure 2), and overall AUC can drop by 20% when testing image classifiers with horizontally flipped images (Figure 3(c)).
+This is expected as the transformed images are, by definition of MIA, non-members.
+To address this gap, we propose a new class of inference attack, Range Membership Inference Attack (RaMIA), which is specifically designed to audit this new type of information leakage (Fig 1).
+The goal of RaMIA is to determine whether any training data exists within a defined neighborhood around a candidate point.
+Instead of relying on point queries that seek an exact match, RaMIA uses range queries that are defined by a center point, a distance measure that captures the semantics of privacy, and a radius that describes the area of interest.
+In practice, an auditor can craft a range query centered on any data record, image, or text.
+By applying a distance function that preserves the sensitive features (e.g., using the $\ell_{2}$ distance on unimportant features or tokens), the auditor can test whether the model has memorized these sensitive features even when the data tested are not exact replicas of any training data.
+In this way, privacy auditors can more accurately and comprehensively assess the privacy risk associated with sensitive information by tailoring the center, range function, and radius to address their specific concerns.
+RaMIA is a flexible framework that supports a wide range of distance measures, enabling it to quantify various privacy risks based on different notions of proximity.
+Moreover, RaMIA aligns more closely with our intuitive understanding of privacy risk: when assessing the privacy of a data point, we expect the test to go beyond detecting exact matches and instead capture if any sensitive information is being memorized.
+For instance, if we claim that a model does not leak private information about my photo, we expect that it does not reveal any private information about me instead of merely the exact test photo.
+The range MIA enables formulating this notion of privacy risk.
+RaMIA is also naturally related to current evaluation protocols for data extraction attacks on generative models [11, 21, 22], where candidate data points that fall within a specific tolerance are treated as successful extractions, reflecting the fact that close-by points can expose private information.
+Moreover, RaMIA can bypass naive test-time privacy protection schemes in which inputs that are recognized as training data are filtered (to fail naive use of MIA).
+Since the model disrupts the correspondence between inputs and exact training points, MIA is not expected to work properly.
+On the other hand, because RaMIA is capable of detecting information leakage in points close to training data, it can easily bypass this simple defense, thereby offering a more robust way of auditing information leakage against data reconstruction attacks.
+Range membership inference attacks extend the formulation of exact membership inference attacks by incorporating range queries that capture privacy leakage in the vicinity of training data.
+We modify the traditional membership inference game to accommodate range queries, leading to composite hypotheses in the likelihood ratio tests, which are the standard and most effective techniques in MIAs [14, 16, 17, 18].
+Our method leverages robust statistical approaches for composite hypothesis testing, including generalized likelihood ratio tests (GLRTs) and Bayes factors.
+We show that RaMIA provides a more comprehensive notion of privacy by detecting leakage from nearby training data when the direct application of MIAs underestimates the true risk.
+For example, testing horizontally flipped images instead of the original training and test images leads to a 20% drop in AUC when audited by MIAs (see Fig 3(c)).
+As a proof-of-concept, we implement RaMIA with a straightforward attack strategy and conduct experiments on tabular, image, and text datasets, where RaMIA consistently outperforms traditional MIA.
+Notably, even in the most challenging scenario where each IN range contains only one training point positioned near the boundary, our simple RaMIA algorithm (Sec IV) achieves a 5% improvement on image datasets (see Figs 4(b) and 4(c)).
+This result is obtained using at most 15 samples, which represents a negligible cost given the high dimensionality of the data space.
+It is important to note that these gains are observed when comparing RaMIA with the case where only traditional MIA is available and the MIA score of the center point is used to audit privacy with RaMIA’s membership definition.
+RaMIA’s performance further improves in less restrictive settings where each IN range contains multiple training points, underscoring its robustness and practical impact for comprehensive privacy auditing.
 (a) MIAScore=2.99
@@
 (c) MIAScore=1.8 and 1.0
-Figure 1: Examples of similar data with drastically different MIA scores.
-MIA would classify one of them as member, but the other as non-member.
+Figure 2: Examples of similar data with drastically different MIA scores and potentially opposite membership predictions.
 The dog image is from the CIFAR-10 dataset, while the text data is from AG News.
-The image classifer is not trained with horizontal flipping augmentation.
-2 Preliminaries
-2.1 Membership inference attacks
-The membership inference attack (MIA) [40] is a type of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set.
+The image classifier is not trained with horizontal flipping augmentation.
+II Related Work
+II-A Membership inference attacks
+The membership inference attack (MIA) [13] is a class of inference attack against machine learning models to infer whether a given data sample is part of the model’s training set.
 Mathematically, given a model $f$ and a query point $\mathbf{x}$, the MIA aims to output 1 if $\mathbf{x}$ is a training point, and 0 otherwise.
-There are various methods to construct and conduct the attack, and it is still an active research direction that sees more powerful attacks being developed.
+Over time, various methods have been developed to construct and execute MIAs, making it an active research area with continuously evolving techniques.
 Shokri et al.
-[40] use a shadow model-based approach where shadow models are trained on known training sets in similar ways to the target model.
-Confidence values of the training and test data on the shadow models are computed, which are then used as benchmarks in testing.
+[13] introduced a shadow model-based approach.
+In this method, multiple shadow models are trained on datasets that mimic the target model’s training set, and the confidence values for both training and test data are computed to serve as benchmarks.
 However, the high cost and strong assumption of knowing the target model’s training details make the attack often infeasible.
-Yeom et al.
 [23] use model loss as a signal and threshold it, scraping the need for shadow models.
-Then MIA is formulated as an inference game (See Sec 3.1.1).
-Researchers turn to the principled approach to solve the game via likelihood ratio tests [39, 5, 45, 48].
-Carlini et al.
-[5] and Ye et al.
-[45] propose reference-model based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio.
-To further boost the attack power, Zarifzadeh et al.
-[48] assumes the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be calibrated on ratios obtained on (non-member) population data.
+Then MIA is formulated as an inference game (See Sec III-A1).
+Subsequent research has adopted a more principled approach by solving this inference game using likelihood ratio tests [14, 17, 16, 18].
+[17] and [16] propose reference model-based approaches, where target signals are compared to those obtained on reference models to obtain the likelihood ratio.
+To further boost the attack power, [18] assumes the attacker has access to a pool of population data so that the likelihood ratio from reference-based attacks can be further calibrated on population data.
 Membership inference attacks with augmentations
 Recent attacks [17, 18] find extra attack performance on image data from augmenting the test queries with train-time augmentations, as these augmented training images could have been seen by the model during training.
-However, this assumes that the attacker knows the exact train-time augmentations in advance, and is able to sample from them.
-Augmenting training images with non train-time augmentations is not considered and allowed for valid reasons: those augmented images would be non-members in the current privacy notion based on membership inference, and testing them as members would violate the current privacy notion.
-2.2 Range queries
-If we make a connection to the field of databases, the existing membership inference attack operates on point queries or exact match queries.
-That is, each query to the membership inference attack only contains one data point and the attack only concerns if this very point is in the training set.
-On the other hand, range query, which is also a common querying operation in database systems, wants to retrieve all data points that fall into the ”range”.
-The most fundamental difference to point query is that the retrieved result often contains multiple data points instead of a single one.
-Our proposed attack, the range membership inference attack, operates with range queries.
-3 From MIA to RaMIA
-Membership inference attacks are often formulated as an inference game [46, 21, 45, 5, 48] between a challenger and an adversary.
-In this section, we will walk through how we come up with RaMIA from MIA.
-3.1 Membership inference attacks
-In membership inference attacks, the goal is to identify if a given point is part of the training set.
-3.1.1 Membership inference game
+This approach, however, presumes that the attacker knows the exact train-time augmentations and can replicate them.
+Augmenting training images with non train-time augmentations is deliberately excluded, as such images would be classified as non-members under the current privacy notion, and testing them as members would contradict the established definition.
+Perturbation-based membership inference attacks
+Several approaches have explored utilizing neighboring points around the query to improve MIA performance.
+[24] hypothesize that membership scores should remain consistently high in the vicinity of training points, and they aggregate scores from neighboring points to assess the smoothness of the scoring function.
+In contrast, [25] argue that the loss curvature would be sharper around training points, causing MIA scores to decline rapidly as one moves away from the training data.
+Although these perturbation-based methods incorporate neighboring information, they ultimately classify only the exact training points as members.
+Any perturbed or neighboring point is treated as non-member.
+II-B Range queries
+Drawing a parallel with database systems, traditional MIAs operate on point queries or exact match queries, where a single data point is retrieved.
+In contrast, a range query is designed to retrieve all data points within a specified interval or “range.” The key difference is that a range query often returns multiple data points rather than a single exact match.
+Our proposed attack, the range membership inference attack (RaMIA), builds upon this concept by operating with range queries, thereby extending the conventional MIA framework to capture privacy leakage in a broader context.
+III From MIA to RaMIA
+Membership inference attacks (MIAs) have traditionally been formulated as an inference game [23, 26, 16, 17, 18] between a challenger and an adversary.
+In this section, we first review the standard MIA framework and its evaluation protocol, then discuss its intrinsic limitations as a privacy auditing tool.
+These limitations motivate our proposed extension, range membership inference attacks (RaMIAs), which broaden the notion of membership to include points that leak sensitive information while not exactly in the training set, .
+III-A Membership inference attacks
+In membership inference attacks, the goal of the attacker is to identify if a given point is part of the training set.
+III-A1 Membership inference game
 Definition 1
@@
   1. 1.
-    The challenger samples a training dataset $D\overset{s_{D}}{\longleftarrow}\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
+    The challenger samples a training dataset $D\longleftarrow\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
   2. 2.
-    The challenger samples a data record $z_{0}\overset{s_{z_{0}}}{\longleftarrow}\pi$ from the data distribution, and a training data record $z_{1}\overset{s_{z_{1}}}{\longleftarrow}D$.
+    The challenger samples a data record $z_{0}\longleftarrow\pi$ from the data distribution, and a training data record $z_{1}\longleftarrow D$.
   3. 3.
@@
     If $\hat{b}=b$, output 1 (success). Otherwise, output 0.
-3.1.2 Evaluation of MIA
-Evaluation is done with a set of training and test points.
-True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all possible threshold values.
+III-A2 Evaluation of MIA
+Conventionally, the MIA algorithm outputs a continuous MIA score for each point query.
+The membership decision is obtained by thresholding the score.
+Evaluation of MIAs is done on a set of training and test points.
+True positive rate (TPR) and false positive rate (FPR) are computed by sweeping over all threshold values.
 By plotting the receiver operating characteristic curve (ROC), the power of an attack strategy can be represented by the area under the curve (AUC).
 A clueless adversary who can only randomly guess the membership labels is expected to get an AUC of $0.5$.
-For stronger adversaries, they predict membership more accurately at each error level.
+Stronger adversaries predict membership more accurately at each error level.
 Hence, they would achieve higher TPR at each FPR, and get a higher AUC.
-3.1.3 Intrinsic limitation of MIA as a Privacy Auditing Framework
-MIAs are intrinsically incapable of identifying points close to training points, regardless of how similar they are, because these points are, by definition, non-members in the scope of MIAs.
-Hence, there is a huge space of points that contain private information but are deemed non-members in the current privacy auditing framework.
-In this way, MIAs as privacy auditing tools become bad when the queries move away from the original data.
-Figure 2 shows the MIAs under-perform on non-original data.
-This inspires our formulation of RaMIA, where these points will be classified as ”members” for better and more comprehensive privacy auditing.
+III-A3 Intrinsic limitation of MIA as a Privacy Auditing Framework
+MIAs are designed to be incapable of identifying points close to training points, regardless of how similar they are, due to the strict definition of members.
+Hence, there is a vast data space of points that contain private information but are deemed to be non-members in the current privacy auditing framework.
+In this way, MIAs as privacy auditing tools become out-of-scope when the queries move away from the original data, resulting in unpredictable and unreliable auditing results in these scenarios.
+Figure 3 illustrates how MIA performance deteriorates as the query points deviate from the original training data.
+This observation motivates our formulation of RaMIA, where we broaden the definition of membership to capture privacy leakage from similar data points.
 (a) Purchase-100
@@
 We define points different from training points but carry similar information as members.
-In 2(a), $m$ is the number of missing values.
+In 3(a), $m$ is the number of missing values.
 The query is constructed by filling in them with the most likely values.
-In 2(b), the point query changes to photos of the same identity who has at least one photo in the training set.
-In 2(c), the queries are horizontally flipped images.
-In 2(d), $d$ is the Hamming distance to original sentences.
-3.2 Range membership inference attack
-In range membership inference attacks, the goal is to identify if a given range contains any training point.
-3.2.1 Range membership inference game
-Here we define our range membership inference game, modified from the above formulation.
+In 3(b), the point query changes to photos of the same identity who has at least one photo in the training set.
+In 3(c), the queries are horizontally flipped images.
+In 3(d), $d$ is the Hamming distance to original sentences.
+III-B Range membership inference attack
+In range membership inference attacks, the goal is to identify if a given range contains any training point. We define our range membership inference game, modified from the MI game.
 Definition 2
@@
   1. 1.
-    The challenger samples a training dataset $D\overset{s_{D}}{\longleftarrow}\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
+    The challenger samples a training dataset $D\longleftarrow\pi$, and trains a model $\theta\longleftarrow\mathcal{T}(D)$.
   2. 2.
-    The challenger samples a data record $z_{0}\overset{s_{z_{0}}}{\longleftarrow}\pi$ from the data distribution, and a training data record $z_{1}\overset{s_{z_{1}}}{\longleftarrow}D$.
+    The challenger samples a data record $z_{0}\longleftarrow\pi$ from the data distribution, and a training data record $z_{1}\longleftarrow D$.
   3. 3.
     The challenger flips a fair coin to get the bit $b\in\{0,1\}$.
-If $b=1$, the challenger samples a range $\mathcal{R}_{1}$ containing at least one training point.
-Otherwise, challenger samples a range $\mathcal{R}_{0}$ containing no training points.
+If $b=1$, the challenger samples a range $\mathcal{R}_{1}$ containing $z_{1}$.
+Otherwise, challenger samples a range $\mathcal{R}_{0}$ containing $z_{0}$ and no training points.
   4. 4.
@@
     If $\hat{b}=b$, output 1 (success). Otherwise, output 0.
-The main difference between the two games is that the adversary now receives a range query (Step 4 in Def 2) instead of a point query (Step 3 in Def 1).
-We assume that the adversary is able to sample a set of points covered by any given range, which effectively means sampling within the range.
-This is a reasonable assumption because the adversary is usually assumed to have the ability to sample from the original data distribution $\pi$ [40, 45, 48] as specified in the membership inference game formulation.
-Given a range query and a sampler of $\pi$, it is not difficult to sample within the range.
+The key difference is that the adversary now receives a range query (Step 4) rather than a single data point.
+We assume that the adversary can sample a set of points from any given range—a reasonable assumption given their ability to sample from the data distribution $\pi$ as in traditional MIAs [13, 16, 18].
 What is a range
@@
 We refer to the center as the query center, the radius as the range size, and the distance function as the range function in this paper.
-One way to visualize a range is to imagine a unit $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions.
-Our framework can cater to any arbitrary range function that is potentially privacy-preserving.
-It can be spatially based (e.g.
-$l_{p}$ distances), transformation based (e.g.
-geometric transformations), and semantic based (e.g.
-owner/main features of the data).
+Formally, we can define a range by =$\mathcal{R}=\{x^{\prime}:d(x^{\prime},x)\leq\epsilon\}$, where $x$ is the range center, $d$ is the range function and $\epsilon$ being the range size.
+One way to visualize a range is to imagine an $l_{2}$ ball around a point $x$, replacing the radius and $l_{2}$ distance with any arbitrary choice of range sizes and functions.
+Our framework is flexible to accommodate any distance function that preserves a significant amount of the sensitive information.
+The range function can be spatial (e.g.
+$l_{p}$ norms), transformation-based (e.g.
+geometric transformations), or semantic (e.g., based on user identity).
 In the experiment section, we will present results with all of these types of range functions.
-Note that our attack reduces to user-level inference [31, 23, 27, 11, 10] when the range function is user-based, which means the privacy concern is whether any other data from the same owner are used in training.
-We will comment on the relations between RaMIA and other existing inference attacks later in the paper.
+Notably, RaMIA reduces to user-level inference [27, 28, 3, 29, 30] when the range function is defined on a per-user basis.
 How to construct a range
-In Step 3 of the range membership inference game, the details of how the challenger samples the ranges are intentionally omitted.
-This is because the ranges can be constructed around either in-distribution or out-of-distribution data points for both IN and OUT-ranges.
-The details of how we construct the iIN and OUT-ranges for our experiments are elaborated in Section 5.1.
-3.3 Evaluation of RaMIA
-Similar to MIAs, we evaluate RaMIA with AUCs.
-However, the notion of true positives and false positives are different from those defined in MIA, as both are defined on the range level.
-Specifically, if a range contains at least one training point, it is IN; otherwise, it is OUT.
-The TPR in RaMIA is the proportion of IN-ranges that are correctly identified by the attacker.
-On the other hand, the FPR in RaMIA is the proportion of OUT-ranges that yield positive predictions by the attacker.
-To avoid confusion, we call them (Range) TPR and (Range) FPR.
-4 Range membership inference attacks
-4.1 (Simple) Hypothesis testing
+In Step 3 of the range membership inference game, the specific procedure for constructing ranges is intentionally left unspecified.
+This is for flexibility: ranges can be constructed around either in-distribution (ID) or out-of-distribution (OOD) data points for both IN and OUT cases.
+The details of our range construction methods for experiments are provided in Section V-A.
+III-C Evaluation of RaMIA
+RaMIA is evaluated similarly to MIAs using AUC metrics, but with definitions adapted to the range setting.
+A range is considered IN if it contains at least one training point and OUT otherwise.
+Thus, the TPR is defined as the proportion of IN ranges correctly identified by the adversary, and the FPR is the proportion of OUT ranges incorrectly classified as IN.
+To avoid confusion, we call them (Range) TPR/FPR.
+IV Range membership inference attacks
+IV-A (Simple) Hypothesis testing
 The standard way to tackle the inference game (Def 1) is to apply statistical hypothesis tests [16, 17]:
@@
 The scoring function in the membership inference attack can be considered as an approximation of the likelihood function $\text{I\kern-1.49994ptP}{(\cdot)}$.
-4.2 Composite hypothesis testing
-Similar to likelihood ratio tests for membership inference games 4.1, we can also construct two hypotheses for range membership inference game (Def 2):
+IV-B Composite hypothesis testing
+When extending to range membership inference, the adversary is presented with a set of points rather than a single point. For a given range $\mathcal{R}_{b}$, we define the hypotheses as:
 […]
@@
 […]
-Since it can be intractable to iterate over all points in a given range $\mathcal{R}_{b}$, we use a set of sampled points $S$ in the given range as its proxy. In this way, the two hypotheses are:
-[…]
-[…]
-[…]
-[…]
+Since it can be intractable to iterate over all points in a given range $\mathcal{R}_{b}$, we use a set of sampled points $S$ in the given range as its proxy to replace $\mathcal{R}_{b}$ in the hypotheses above.
 The likelihood ratio in this case is still $\frac{\text{I\kern-1.04996ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.04996ptP}{\left(\theta|H_{0}\right)}}$.
-However, the alternative hypothesis $H_{1}$ is composite because it is a union of multiple hypotheses $\bigcup_{z_{i}\leftarrow S}(z_{i}\in D)$.
+However, the alternative hypothesis $H_{1}$ is composite, representing a union of multiple hypotheses $\bigcup_{z_{i}\leftarrow S}(z_{i}\in D)$.
 Therefore, we need to use statistical methods tailored for composite hypothesis testing.
-There are two commonly used methods for it: Bayes Factor [22] and Generalized Likelihood Ratio Tests (GLRTs) [43].
+Two commonly used methods are Bayes Factor [31] and Generalized Likelihood Ratio Tests (GLRTs) [32].
 GLRT
-GLRT assumes that the true hypothesis $h^{*}$ is explicitly listed in the composite hypothesis, which, in our case, means the training point is in the set of samples $\{z_{i}\}$.
-In this case, the composite hypothesis can be reduced to a single hypothesis by taking the maximum:
+GLRT assumes that the true hypothesis $h^{*}$ is explicitly present in the composite hypothesis, which, in our case, means the training point is in $S$.
+This allows us to reduce the composite hypothesis by taking the maximum over $S$:
 $\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}\Rightarrow\max_{x\in S}\text{I\kern-1.49994ptP}{(\theta|x\in D)}.$ (2)
 Bayes Factor
-Bayes Factor assumes the hypothesis is a random variable, each with a likelihood value according to a prior distribution.
-In our case, it means each point $z$ in the range can be sampled with probability determined by the prior data distribution $\text{I\kern-1.49994ptP}{(z)}$.
-To reduce the composite hypothesis to a simple one, Bayes Factor computes the weighted average:
+Bayes Factor treats the hypothesis as a random variable with a prior distribution.
+Each point z in the range is sampled with probability determined by $\text{I\kern-1.49994ptP}{(z)}$, and the composite likelihood is approximated via the expectation:
 […]
-This can be thought of as taking the expectation of the composite hypothesis or using the most representative hypothesis instead of the maximal (as in GLRT) in the likelihood ratio test.
 Why both methods fall short in RaMIA
-At first look, GLRT might be an intuitive choice, because taking the maximum leads to a two-step strategy: search and test.
+At first look, GLRT might be an intuitive choice, as it is equivalent to a two-step strategy: search and test.
 Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points, and their membership should be indicative of the ranges’ membership.
-However, this assumes the true training point is included in every sampled set created by the adversary, which is extremely unlikely.
+However, this assumes the true training point is in every sampled set $S$, which is extremely unlikely.
 Secondly, this also assumes that we can reliably find the max values in a given range.
-Since most ranges are large data subspaces, it is very challenging to find the optimal points within the large space.
-Even if the search space can be navigated, any search algorithm is likely to return local maxima.
-To make full use of the Bayes Factor, we need to know the prior distribution, which is also not possible.
-A more important issue is that the power of existing statistical methods relies on the absolute correctness of the likelihood values: the larger the likelihood values, the more likely the point is a member.
-In other words, the largest likelihood values of non-members must be smaller than the smallest likelihood of members.
-However, membership inference attacks are known to be non-perfect; they are also known to be particularly unreliable on out-of-distribution (OOD) data [48] and can assign them high scores.
-When the sampling space contains only these data as opposed to real and in-distribution (ID) data, the maximum might not be anything close to true training data, increasing the FPR and lowering the AUC as a result.
-4.3 Our approach: Trimmed averages
-Given these limitations, in this paper, we propose a simple attack strategy based on Bayes Factor.
-We first assume every sampled point is of equal probability.
-Then, to address the unreliability of MIA scores, we use a trimming process based on the type of data in the sampling space.
-If the data are all naturally ID, we trim the less probable samples and take the average likelihood of the top samples to reduce the influence of non-members and the randomness in the sampling process.
-On the other hand, if the adversary can only synthesize data within the range instead of sampling from a pool of real data, the top samples are highly likely to be OOD data with high membership scores.
-Hence, in this case, we want to remove those points from the equation.
-We then trim the top-scoring samples and take the average of the rest.
-This gives us the following:
+Since most ranges are large data subspaces, it is very challenging to find the extreme points.
+Even if the search space can be navigated, search algorithms are likely to return local maxima.
+Similarly, the Bayes Factor approach requires knowledge of the true prior $\text{I\kern-1.49994ptP}{(x)}$, which is generally unknown.
+Both methods also hinge on the absolute correctness of the likelihood values: ideally, the highest likelihood among non-members should be lower than the lowest likelihood among members.
+However, membership inference attacks are known to be imperfect and particularly unreliable on OOD data [18], which may assign high scores erroneously to non-members.
+This leads to increased Range FPR and reduced AUC, especially when working with sampled data.
+IV-C Our approach: Trimmed averages
+To overcome these limitations, we propose a robust attack strategy based on a modified Bayes Factor approach that employs trimmed averages.
+Our method begins by assuming every sampled point in $S$ is equally probable.
+To mitigate the unreliability of MIA scores, we introduce a trimming process that adapts to the nature of the data in the sampling space:
+  • In-Distribution (ID) Data: When the sampled points are naturally in-distribution, we trim the lower quantiles (i.e., those with the smallest likelihood values) and average the top samples.
+This reduces the influence of non-members and random noise.
+  • Synthetic Data: When only synthetic data are available, the highest scoring samples are often OOD and prone to false positives.
+In this case, we trim the top quantiles and average the remaining samples.
+This strategy is formalized as:
 $\end{multlined}\text{I\kern-1.49994ptP}{\left(\theta|H_{1}\right)}=\text{TrimmedAvg}(S,q_{s},q_{e};\text{I\kern-1.49994ptP})\\ =\text{Avg}_{x\not\in[q_{s},q_{e}]\text{-th quantiles}}\text{I\kern-1.49994ptP}{(\theta|x\in D)},$ (4)
-where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles that we want to remove to compute our robust statistics that are one-sided trimmed means.
-If the sampling space is filled with synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points.
+where $S$ is the sampled set, $q_{s}$ and $q_{e}$ denote the quantile thresholds between which to be trimmed.
+For synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points.
 $q_{s}$ is a hyperparameter that decreases (trim more) as the quality of sampled points gets worse.
-On the other hand, if the sampling space consists of real points, we set $q_{s}=0$ to remove the smallest points in our aggregation.
+For real data, we set $q_{s}=0$ to remove the smallest points in our aggregation.
 $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available.
 If we do not trim anything, the formulation reduces to Bayes Factor which assigns equal probability to all points.
-Note that the optimal hyperparameters may differ across different membership signals (e.g.
-loss values, LiRA scores), as they exploit different vulnerabilities and expose different training points.
-However, for fixed model architectures, range functions, data distributions, and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [48].
-Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters in a grid search manner.
-More details are in Appendix A.
-4.4 Range membership inference attack as a framework
-The range membership inference attack is a new inference attack framework, not a particular attack algorithm.
-There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 4.
-The sampler $\text{Sampler}(\mathcal{R}):S\rightarrow X$ returns samples within the given range.
-The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\text{I\kern-1.49994ptP}{(\theta|x)}$.
-Any existing MIA algorithm can be plugged in as RaMIA’s membership testing backbone.
-Similar to MIAs, the key to using RaMIA as a privacy auditing tool is to compute the range membership score.
-Our framework can adopt any existing membership scoring function $\text{MIA}(x)$ to compute $\text{RaMIA}(\mathcal{R})$.
-Below, we outline the attack with our attack strategy described above:
+We provide an empirical comparison in Fig 11 to show that the trimmed means are better suited for RaMIA.
+Note that the optimal hyperparameters may vary with different membership signals (e.g., loss values, LiRA scores) since these signals capture different vulnerabilities.
+For fixed model architectures, range functions, data distributions, and sampling methods, these hyperparameters can be determined using reference models, similar to the offline version of RMIA [18].
+By randomly choosing a reference model as a temporary target model and sweeping through hyperparameters via grid search, one can identify the best values (details in Appendix -A).
+IV-D Range membership inference attack as a framework
+Our proposed RaMIA is not a single attack algorithm but a new inference attack framework. It comprises two key components:
+  • Sampler: $\text{Sampler}(\mathcal{R}):\mathcal{R}\rightarrow S$ returns samples within the given range
+  • Membership Tester: MIA(x) is any point-query membership inference algorithm that outputs a membership score, approximating $\text{I\kern-1.49994ptP}{(\theta\mid x)}$.
+This function can be replaced by any existing MIA algorithm.
+The core idea is to compute a range membership score $\text{RaMIA}(\mathcal{R})$ by robustly aggregating the membership scores of the sampled points $S$ using our trimmed average approach (Eqn.
+4).
+This framework leverages existing membership scoring functions while addressing the new challenge of capturing privacy leakage in a broader neighborhood.
 Algorithm 1 Computing range membership scores
@@
 8:$\text{RaMIA}(\mathcal{R};\theta)=\text{TrimmedAvg}(S,q_{s},q_{e};\text{MIA})$
-5 Experiments
+Importantly, our formulation of RaMIA as a composite hypothesis test—and the subsequent development of the trimmed averages approach—represents a first, yet promising, attempt to tackle this new technical challenge.
+While there is room for further exploration and refinement, our approach demonstrates that extending the membership inference framework to encompass ranges can capture significant privacy leakage in the vicinity of training data.
+This represents a core contribution of our work and lays the groundwork for future research in comprehensive privacy auditing.
+V Experiments
 Since the purpose of this paper is to introduce a new concept and framework, the goal of the experiments section is to provide a proof-of-concept.
 We experiment on the commonly used Purchase-100 [13], CelebA [33], CIFAR-10 [34] and AG News [35] datasets.
-We will describe the details of how we split the dataset, train models, construct ranges, and obtain samples in Section 5.1.
+Details on dataset splits, model training, range construction, and sample acquisition are described in Section V-A.
 Since the range membership notion is new, we do not have a prior method to compare with.
-However, since the goal of introducing this new privacy notion is to enable better and more comprehensive privacy auditing, we compare RaMIA with the de-factor privacy auditing framework, MIA.
-We show that RaMIA performs better (Sec 5.5) in scenarios where MIA under-performs (Depicted in Figure 2).
-Both MIA and RaMIA in the experiments are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [48], with a few reference models trained in the same way as [5] and [48].
-The respective queries are outlined in Table 1, and the definitions of members under each attack framework are explained in Table 2.
-In both tables, $x$ represents original data in datasets, while $x^{\prime}$s are either data with missing values or modified data from $x$.
-The reason that we do not test our attacks by taking ranges centered at original data $x$ is that the chance of the attack data being exactly the same as the training data is extremely low without sufficient prior knowledge.
-It is more realistic that similar data are being queried.
-Table 1: Range queries and point queries used in our experiments for RaMIA and MIA respectively.
+However, as our aim in introducing this new privacy notion is to enable better and more comprehensive privacy auditing, we compare our RaMIA framework with the de facto privacy auditing framework, MIA.
+Note that It is crucial to emphasize that our experimental comparison between RaMIA and MIA is not based on the standard MIA notion of membership.
+In standard MIAs, only the exact training data are considered members.
+In our experiments, the queries for MIA are the range centers—which are not training data by definition.
+Consequently, if one were to evaluate MIA using the correct MIA privacy notion, the AUC would be close to random guessing (i.e., 0.5).
+Instead, we use the MIAScore of the range center to solve the range membership inference game, even though this deviates from the standard definition.
+This allows us to compare the power of both frameworks in identifying queries that leak privacy.
+Table I outlines the range queries and point queries used for RaMIA and MIA respectively, while Table II defines the notion of members under each attack setting.
+In both tables, $x$ represents the original data in the datasets, while $x^{\prime}$ represents either data with missing values or modified data derived from $x$.
+The reason we do not center ranges at the original data $x$ is that, without sufficient prior knowledge, the probability of the attack data exactly matching a training point is extremely low.
+In practice, similar but not identical data are more likely to be queried.
+It also acts as a hard case for RaMIA.
+TABLE I: Range queries and point queries used in our experiments for RaMIA and MIA respectively.
 Dataset Range query Point query
@@
 AG News sentences that are of Hamming distance 8 to sentence $x^{\prime}$ sentence $x^{\prime}$
-Table 2: Definitions of range and point members corresponding to the attack queries in Table 1.
+TABLE II: Definitions of range and point members corresponding to the attack queries in Table I.
 Dataset Range member if there is at least (Point) member if
@@
 AG News one training sentence within Hamming distance 8 to $x^{\prime}$ $x^{\prime}$ is member
-5.1 Setup details
-As mentioned earlier, the range function must come from the semantics of privacy.
-Hence, in experiments, we use specific range functions with different data types.
-For tabular data, we consider missing columns, which is an extreme case of using Euclidean distance on missing columns as a range function.
-For human photo data, we use a semantic range function, which is based on the main person featured in the photo.
+V-A Setup details
+As mentioned earlier, the range function must derive from the semantics of privacy.
+Hence, in experiments, we use specific range functions tailored to different data types.
+For tabular data, we consider missing columns, which is an extreme case of using Euclidean distance on missing columns as a range function ($R=\{x^{\prime}:d(x^{\prime}_{i},x_{i})=0\land d(x^{\prime}_{j},x_{j})<C_{j}\}$, where $j$’s are masked columns and $i$’s are observable columns, $d$ is Euclidean distance, $C$ is an upper bound for each missing column to make sure $x^{\prime}_{j}$ is bounded by the infinity or extreme values (e.g., Age.).
+For human photos, we use a semantic range function based on the main person featured.
 For other image data, we use geometric transformations as range functions.
-For text data, we use (word level) Hamming distance, which is edit distance or Levenshtein distance that only considers word substitution.
+For text data, we use (word level) Hamming distance, which is edit or Levenshtein distance that only considers word substitution.
 The reasons for choosing these range functions have been motivated and explained in earlier parts of this paper.
-5.1.1 Tabular data: Purchase-100
-  • Dataset Purchase-100 [40] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data.
-This dataset was first curated by [40] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product.
-The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
-  • Models We train five-layer multi-layer perceptron (MLP) models in PyTorch [34] on half of the entire dataset.
-The hidden layers are of sizes $[512,256,128,64]$, identical to the setup in [48].
-All models achieve a test accuracy of $83\%$.
+V-A1 Tabular data: Purchase-100
+  • Dataset Purchase-100 [13] is derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data.
+It contains 600 binary features, representing the purchase history of each person.
+The data is divided into 100 classes.
+The task is to predict a person’s category given the purchase history.
+  • Models We train a four-layer multi-layer perceptron (MLP) in PyTorch [36] on half of the entire dataset.
+The hidden layers are of sizes $[1024,512,256]$.
+All models achieve a test accuracy of $86\%$.
   • Construction of ranges We simulate the scenario where the attacker has incomplete data (data with missing values).
 For all training and test data records, we randomly mask k columns.
-Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary.
-We then check if each range constructed from test data points includes any training point.
-If so, they are re-labeled as ”IN-ranges”.
+Each row with masked columns is a range query that contains $2^{k}$ possible points.
+We re-labeled any range that contains at least one training point as ”IN”.
   • Sampling within ranges Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns.
 The parameter of the sampler is computed by taking the average value of each column.
-Because of the nature of this dataset, our sampled data can be regarded as in-distribution.
+Due to nature of this dataset, our sampled data can be regarded as ID.
 We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
-5.1.2 Image data I: CelebA
+V-A2 Image data I: CelebA
   • Dataset CelebA [33], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features.
-We construct the members set by only including photos of celebrities with identity numbers smaller than 5090.
+We construct the members set by only including photos of the first 5090 celebrities.
 The rest are used to construct the non-members set.
@@
 Our target model has a test accuracy of $87\%$.
-  • Construction of ranges The range function here is a semantic one that is based on the identity of the face image.
-For example, a range query can be ”all Alice’s photos”.
-Since the identities in the training and non-members set are disjoint, it is easy to construct IN- and OUT-ranges based on the distribution of identities in the two sets.
-  • Sampling within ranges For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
-5.1.3 Image data II: CIFAR-10
-  • Dataset CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
+  • Construction of ranges The range function here is semantic, defined by identity.
+For example, a range query can be ”all photos of Alice”.
+Since the identities in the training and non-members set are disjoint, it is easy to construct IN and OUT ranges.
+  • Sampling within ranges For each range query, we collect all holdout images sharing the same identity as the range center.
+V-A3 Image data II: CIFAR-10
+  • Dataset CIFAR-10 [34] is a popular image classification dataset, containing 50,000 training images of size $(32,32,3)$.
   • Models We train WideResNets-28-2 [37] with JAX [38] on half of the training set of CIFAR-10 using the code from [17], with and without image augmentations.
-Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$.
-The train time augmentation is the composition of random flipping, cropping and random hue.
-  • Construction of ranges The range function here is different types of image augmentations, which are geometric transformations.
-An example of a range query is ”all transformed version of image $X$”.
-For each training and test image, a range is constructed by applying different transformations.
-  • Sampling within ranges For each range query, we independently apply 15 image augmentations on the query center.
+Our target model achieves a test accuracy of $83\%$ when trained without augmentation, and $92\%$ with augmentation.
+The train-time augmentation is the composition of random flipping, cropping and random hue.
+  • Construction of ranges The range function is defined by a set of geometric transformations (e.g., flipping, rotation, cropping).
+A range query consists of various transformed versions of image $x^{\prime}$.
+  • Sampling within ranges For each range query, we independently apply 10 image augmentations to the range center.
 The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
@@
 (d) AG News (d=5)
-Figure 3: RaMIAs perform better than MIAs on points that are close to original points but not exactly the same.
-5.1.4 Textual data: AG News
-  • Dataset We use the AG News dataset [49], which is a news collection with four categories of news.
-Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset.
-There are 120,000 sentences in its training set.
-It is a popular text dataset in LLM MIA literature [33].
-  • Models We took pretrained GPT-2 [35] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library.
+Figure 4: RaMIAs perform better than MIAs on similar but not identical data points to training points that carry overlapping private information.
+This is the hardest setting for RaMIA where there is only one training point in each IN range, and they are not situtated at the range centers.
+We will show how RaMIAs will perform better in less restricted scenarios in a later figure.
+Note that the TPRs/FPRs are the range versions, and the MIA is the approach that uses the range center’s MIA score as the range’s membership signal, both introduced in Sec V-D.
+V-A4 Textual data: AG News
+  • Dataset We use the popular AG News dataset [35], which is a news collection with four categories of news.
+We treat it as a text generation dataset, disregarding their labels.
+It contains 120,000 sentences in the training set.
+  • Models We took pretrained GPT-2 [39] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [40] (implemented in Hugging Face’s PEFT [41] library).
 The finetuning is done for 4 epochs.
-Our target model achieves a perplexity of 1.39 on the test set of AG News.
+Our target model achieves a test perplexity of 1.39.
   • Construction of ranges The range function here is word-level Hamming distance, which can be thought of as the edit distance measured on word level that only allows word substitution.
 An example of a range query is ”all sentences within Hamming distance $d$ to sentence $x$”.
-To construct IN- and OUT-ranges, we just need to specify the max Hamming distance and the starting sentence.
-We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pretrained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences.
-A Hamming distance is then specified with each starting sentence to form a range.
-  • Sampling within ranges We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
-5.2 Hyperparameters
+To construct IN and OUT ranges, we construct range centers by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pretrained BERT [42] model, so they have a distance of $\alpha$ to the original training/test sentences.
+  • Sampling within ranges We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [42] to complete the masks.
+V-B Hyperparameters
 Overall, on Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$.
-On CIFAR-10, we apply up to 15 distinct transforms, and set $q_{e}=100,q_{s}=40$.
+On CIFAR-10, we apply up to 10 distinct transforms, and set $q_{e}=100,q_{s}=40$.
 On AG News, we construct 50 sentences within each range, and set $q_{e}=100,q_{s}=20$.
@@
 We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
-5.3 Implementation details
-We train 8 models on Purchase-100 and CIFAR-10, and train 4 models on CelebA and AG News.
-Each model is trained on half of the dataset, as described in [5, 48].
-For all PyTorch models, we use Adam as our optimizer with a learning rate of 0.001.
+V-C Implementation details
+We train 16 models on Purchase-100, and 4 models on CelebA, CIFAR-10, and AG News.
+Each model is trained on half of its respective dataset, following setups from [17, 18].
+For all PyTorch models, we use Adam with a learning rate of 0.001.
 For WideResNets, we use the training code from [17].
-On AG News, the models are trained for 4 epochs.
-On other datasets, they are trained for 100 epochs.
-All training is done on two Nvidia RTX 3090 GPUs.
-Training on AG News takes about 1 hour per epoch.
-Training other models takes less than one hour each.
-5.4 Metrics
-To evaluate the performance of RaMIA and MIA based on the same MIA algorithm, we use AUCs.
-The inputs to RaMIA are range queries in Table 1, while the inputs to MIA are point queries, which are range centers, in Table 1.
-The ground truth membership for both RaMIA and MIA is the range membership information defined in Table 2, which allows us to compare the power of both inference attack frameworks in identifying queries that are privacy leaking.
-In other words, both point and range TPRs/FPRs are calculated against the range membership information.
-Following [5], we also report (Range) TPR at small (Range) FPR for both MIA and RaMIA (Table 4).
-5.5 RaMIAs quantify privacy risks more comprehensively than MIAs
+AG News models are trained for 4 epochs, and other datasets for 100 epochs.
+Training is conducted on two Nvidia RTX 3090 GPUs, with AG News taking about 1 hour per epoch and other models less than one hour each.
+V-D Metrics
+To evaluate the performance of RaMIA and MIA using the same membership inference backbone, we measure AUCs.
+The inputs to RaMIA are range queries (as specified in Table I), while the inputs to MIA are the point queries corresponding to range centers.
+Importantly, the ground truth membership for both RaMIA and MIA is defined according to range membership (see Table II).
+This means that even though the MIA attack is based on point queries, the true membership is based on whether the range contains at least one training point.
+If evaluated against the correct MIA privacy notion (which only considers exact training points as members), the AUC would be close to random guessing ($\approx 0.5$).
+Our experimental setup uses the MIAScore of the range center as a proxy to solve the range membership inference game, allowing us to compare the power of both frameworks in identifying privacy leaking queries.
+It also illustrates the utility of sampling within the range to construct a more meaningful score under RaMIA’s privacy notion.
+Following [17], we also report (Range) TPR at small (Range) FPR for both methods (see Table IV).
+V-E RaMIAs quantify privacy risks more comprehensively than MIAs
 As we have explained before, data points that are close enough to the training data are out of the scope of membership inference attacks.
-We observe from Figure 3 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested.
-We want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes.
+We observe from Figure 4 and 10 that range membership inference attacks are better at identifying those nearby points, and thus providing more comprehensive privacy auditing on all the four datasets we tested.
+We want to highlight that this is the hardest setting for RaMIA, where each IN range only contains one training point that is also not the range center.
+We also want to emphasize that the gain is remarkable if we consider how little samples were taken compared to the range sizes.
 On Purchase-100, there are a total of 1024 candidates, and we take less than 20% of them.
@@
 Yet, limited samples can lead to noticeable gains, which further shows the current privacy quantification approach is suboptimal and needs a better framework.
-Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table 3.TPRs at small FPRs are in Table 4.
-The improvement in AUC is summarized in Table 3.
-Table 3: Improvement in AUCs after switching from MIA to RaMIA across multiple iterations of random sampling. We do not randomly sample but use all available attack images in CelebA.
+Due to randomness in sampling, we report the average gain of RaMIA over MIA with standard deviation in Table III.TPRs at small FPRs are in Table IV.
+The improvement in AUC is summarized in Table III.
+TABLE III: Improvement in AUCs after switching from MIA to RaMIA across multiple iterations of random sampling. We do not randomly sample but use all available attack images in CelebA.
 Purchase-100 CIFAR-10 CelebA AG News
-$\Delta\text{AUC}$ $2.62\pm 0.04$ $4.12\pm 0.06$ $5.4$ $1.20\pm 0.2$
+[…]
+[…]
 Relation to user-level inference
@@
 In terms of algorithms, our attack strategy should dominate the simple averaging approach used in prior work [27, 28], since the trimming ratio is optimized (compared to no trimming in simple averaging).
-There are other user-level inference algorithms [11] that train additional shadow models and classifiers, which incurs additional computational costs.
-They also consider scenarios where user information is the label of their target models, e.g.
+The comparison can be found in Fig 11(b).
+There are other user-level inference algorithms [29] that train shadow models, which incurs additional computational costs.
+But these methods usually consider scenarios where user information is the label, e.g.
 facial and speech recognition systems that predict user ID as their outputs [29, 30].
@@
 On the other hand, our target model, a facial attribute classifier, does not use user information anywhere in the training, making the inference harder.
-Table 4: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$.
+TABLE IV: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$.
 MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them.
 Standard deviation over random sampling iterations is reported, except for CelebA, where we use all available candidates.
-The TPR and FPR are calculated based on the range membership information, as described in Sec 5.4.
+The TPR and FPR are calculated based on the range membership information, as described in Sec V-D.
 For a fair comparison, we should compare RaMIA and MIA based on the same membership testing backbone, e.g.
 MIA with RMIA versus RaMIA with RMIA.
+As we have argued, the test queries are unlikely to be exact matches to training points, hence the MIAs are evalauted on range centers, which should be considered non-members.
+Hence, the TPR values are not indicative of the attack power.
+This table is only for interested readers who wants to know the attack performance at small FPRs.
 Purchase-100 CIFAR-10 CelebA AG News
@@
 MIA
-LOSS 0 0 0.15 0 1.86 0.31 $80.0$
-RMIA 2.18 0.37 2.40 0.21 1.69 0.19 $0.30$ 00
+LOSS 0 0 0.92 0.02 1.86 0.31 $80.0$
+RMIA 2.18 0.37 0.99 0.09 1.69 0.19 $0.67$ $0.04$
+LiRA 5.20 0.02 2.12 0.57 1.68 0.24 0.68 0.00
 RaMIA
-LOSS $0\pm 0$ $0\pm 0$ $1.17\pm 0.06$ $0.05\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
-RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $3.59\pm 0.11$ $0.80\pm 0.02$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
-5.6 Factors affecting RaMIA performance
+LOSS $0\pm 0$ $0\pm 0$ $0.88\pm 0.06$ $0.09\pm 0.03$ 1.40 0.28 $1.10\pm 0.11$ $0\pm 0$
+RMIA $2.57\pm 1.58$ $0.57\pm 0.47$ $1.41\pm 0.00$ $0.24\pm 0.00$ 1.44 0.22 $0.54\pm 0.12$ $0\pm 0$
+LiRA $2.47\pm 1.47$ $0.50\pm 0.41$ $1.53\pm 0.00$ $0.58\pm 0.00$ 1.10 0.02 $0.63\pm 0.13$ $0\pm 0$
+V-F Factors affecting RaMIA performance
 Training data density in the range
-Due to the nature of the sampling-based approach, the chance of our attack set containing a true training point scales linearly with the density of training points in the range.
-If we keep the sample size constant, increasing the range without including more training points in the range hurts the attack performance because the chance of the attack set including any training point gets diluted.
-On the other hand, if we increase the training point density, which is equivalent to increasing the probability the attacker samples a true training point, the attack performance gets boosted.
+Due to the nature of the sampling-based approach, the chance of our attack set including a true training point scales linearly with the density of training points in the range.
+For a fixed sample size, increasing the range without introducing more training points hurts the attack performance because the chance of the attack set including any training point gets diluted, and vice versa.
 Figure 5(b) shows that the performance of RaMIA increases when the range becomes larger in the CIFAR-10 experiment.
-Recall that the range function in CIFAR-10 is based on image augmentation methods.
-Increasing the range means the attacker applies more distinct augmentation methods to obtain transformed images.
-This increases the chance of the attacker obtaining one of the transformed versions of training images seen by the model during training, thus leading to better attack performance.
-In Figure 3(b), we conducted the attack assuming the attacker cannot sample any true training images.
-As a sanity check, we relax this assumption, and Figure 4(a) shows that RaMIA performs monotonically better when the density of training images increases from 0% to 50%, when the number of samples is constant.
-(a) RaMIA on CelebA gets better when the training points available for sampling increases.
-The percentages are the density of training points in the distribution that the attack samples from.
-(Sanity Check)
-(b) RaMIA performs better on CIFAR-10 when the range size increases. The range size here is equal to the number of distinct transformations applied to images.
+Since the range function in CIFAR-10 is based on image augmentation methods, increasing the range means the attacker applies more distinct augmentation methods, which effectively increases the chance of the attacker obtaining one of the transformed versions of training images seen by the model during training, thus leading to better attack performance.
+In Figure 4(b), we conduct the attack assuming the attacker cannot sample any true training images.
+As a sanity check, we relax this assumption, and Figure 5(a) shows that RaMIA performs monotonically better when the density of training images increases from 0% to 50%, when the number of samples is fixed.
+(a) RaMIA on CelebA gets better when the training points available for sampling increases. The percentages are the density of training points in the sampling space.
+(b) RaMIA performs better on CIFAR-10 when the range size increases. The size is equal to the number of distinct transformations ($n$) applied to images.
 Figure 5: Attack performance increases in tandem with training point density.
+(a) Purchase-100 with $\ell_{1}$ range functions
+(b) CIFAR-10 with $\ell_{2}$ range functions
+Figure 6: Attack performance decreases when increasing range sizes $d$ without introducing more training members to ranges, which effectively reduces the training data density in the range.
 (a) Purchase-100
@@
 It shows that the vulnerability to MIA is positively correlated with that to RaMIA.
+We also observe a monotonic decrease in RaMIA performance as the range size increases without introducing true members (Fig 6).
+We use L1 distance as the range function for Purchase-100, and L2 distance for CIFAR-10, and construct ranges around each data point in MIA’s evaluation set.
+We ensure that each IN range only contains one training point, which is the range center, and that each OUT range does not contain any training point.
 Susceptibility to MIAs and RaMIAs is correlated
@@
 Generally speaking, identities that have more training photos are more prone to RaMIAs.
-Similarly, correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 5.
+Similarly, correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles (Fig 7).
 Figure 8: Number of photos from the same celebrity in the training set affects the identifiability of the range.
-5.7 Mismatched training and attack data hurts attack performance
+V-G Mismatched training and attack data hurts attack performance
 Figure 3(c) shows that MIA underestimates the privacy risk when the augmentation used in training and attacking differs.
-This rings a bell as many people audit the privacy risk of image classifiers with original images, when the classifiers are often trained with a composition of augmentations.
+This should be alarming as many people audit the privacy risk of image classifiers with original images, when the classifiers are often trained with a composition of augmentations.
 Many transformations, such as color jittering and affine transformations, always produce different final images.
@@
 Therefore, we should use RaMIA for a better auditing result (Figure 4(c)).
-Difference to existing augmentation-based MIAs
-Existing attacks [5, 48] also use augmented queries in the attack, but with a different rationale and assumption of the attacker’s knowledge.
-Since they adopt the existing privacy notion based on point queries, only (augmented) images seen by the model in the training stage are considered as members.
-Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy notion.
-In RaMIA, the set of augmentations is given by the challenger (Def 2), which can contain augmentations not used in training, but considered as privacy leaking.
-Using the aggregation method in [5] will hurt the attack performance if non-training augmentations are used.
-However, RaMIA is designed to be robust in this scenario (Fig 3(c)) due to its trimming process.
-5.8 RaMIA on redacted data
+V-H RaMIA on redacted data
 Many large language models (LLMs) are trained with sensitive textual data.
@@
 Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM.
-Accurately identifying the redacted sentences paves the way for reconstructing them as a follow-up attack.
+Accurately identifying the redacted sentences paves the way for potentially better data extraction and reconstruction attacks.
 Figure 9 shows the results.
@@
 Figure 9: RaMIA on a subset of AG News where names are redacted.
-6 Conclusion
-In this paper, we argue that membership inference attacks are only useful as a privacy audit tool when querying exact copies of training and test data.
-Moving the query to similar points causes a drastic decrease in performance, rendering MIAs less useful.
-We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs and covers the failure cases of MIAs by checking if a given range contains a training point.
-We introduce RaMIA as an attack framework that can be implemented with any existing MIA algorithm.
-We show that it can provide better privacy auditing with very few samples taken randomly.
-We hope our work can make more privacy researchers and practitioners aware of the shortcomings of MIAs, and shift their attention to RaMIAs.
-As it is the first paper that brings up this new framework, there is room for improvement in specific attack algorithms.
-For example, a better sampling process will surely increase the gap between RaMIA and MIA.
-Nevertheless, we have shown our framework is sensible and useful.
-In future work, we hope to design more powerful RaMIA strategies that are robust to the change of membership signals and datasets, especially on LLMs where we believe our privacy notion is extremely relevant.
-Appendix A Attack algorithms
-In this section, we explain the details of the membership inference attack algorithms used in our experiments.
-LOSS
-LOSS [46] computes loss values as a proxy of membership score on given points: $\text{MIA}(x;\theta)=\mathit{l}(x;\theta)$.
-To compute the likelihood, an easy way is to take the exponential of the negative of the loss $\text{I\kern-1.49994ptP}=\exp^{-\mathit{l}}$.
-RMIA
-RMIA [48] computes membership score by applying chain rule: $\text{I\kern-1.49994ptP}{(\theta|x)}=\frac{\text{I\kern-1.04996ptP}{(x|\theta)}\text{I\kern-1.04996ptP}{(\theta)}}{\text{I\kern-1.04996ptP}{(x)}}$.
-The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\text{I\kern-1.49994ptP}_{z\in Z}{[\frac{\text{I\kern-1.04996ptP}{(\theta|x)}}{\text{I\kern-1.04996ptP}{(\theta|z)}}\geq\gamma]}$, where the term $\text{I\kern-1.49994ptP}{(\theta)}$ will cancel out with each other.
-The normalizing constant $\text{I\kern-1.49994ptP}{P(x)}$ is computed with reference models: $\text{I\kern-1.49994ptP}{(x)}=0.5\mathbb{E}_{\theta_{\text{IN}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{IN}})}+0.5\mathbb{E}_{\theta_{\text{OUT}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{OUT}})}$.
-In its offline version, the in models are unavailable.
-In this case, the former probabilities are approximated by the latter term $\text{I\kern-1.49994ptP}_{\text{IN}}=a\text{I\kern-1.49994ptP}_{\text{OUT}}+(1-a)$.
-The hyperparameter $\alpha$ is chosen based on the reference models.
-Specifically, one reference model is chosen as the temporary target model, and the rest are used to attack it.
-The value of $\alpha$ is chosen to be the best-performing value under this setting, obtained via a simple sweeping.
-In our experiment, we use the offline attack only.
-The $\alpha$ values for Purchase-100 and CIFAR-10 are taken from [48].
-For CelebA, we set it to be 0.33.
-For AG News, we set it to be 1.0.
+VI Conclusion
+We have shown that traditional membership inference attacks (MIAs), which only consider information leakage at exact training points, fail to capture the broader privacy risks in similar but not identical data.
+We show that when the query shifts from exact training points to nearby points, MIA performance degrades drastically under the correct privacy notion.
+To address this shortcoming, we introduced range membership inference attacks that evaluates whether a given range contains any training data.
+Our formulation casts the problem as a composite hypothesis test and proposes a robust trimmed averaging approach to aggregate membership scores over a set of sampled points.
+Our proof-of-concept experiments on tabular, image and text datasets demonstrate that RaMIA outperforms MIAs in realistic scenarios with non-exact queries.
+We acknowledge that RaMIA presents a new technical challenge and that our current formulation and solution are not optimal; rather, they are intended as a starting point to stimulate further research in this promising and meaningful direction.
+While there is ample room for improving the sampling process and refining the range function design, our work highlights the potential of range-based privacy auditing and motivates future efforts to develop more powerful and robust RaMIA strategies.
+We hope that our work can encourage privacy researchers and practitioners to re-examine the conventional MIA paradigm and consider range-based approaches as a more realistic and comprehensive tool for privacy auditing.
+Acknowledgement
+The authors would like to thank Milad Nasr for the discussion and exploration at the early stage of this project.
+This research is supported by the Ministry of Education, Singapore, Academic Research Fund (AcRF) Tier 1 (A-8001610-00-00) and Tier 2 (MOE-T2EP20223-0015).